### v5

# Surgical Aggregation: Federated Class-Heterogeneous Learning

Abstract
The release of numerous chest x-ray datasets has spearheaded the development of deep learning models with expert-level performance. However, they have limited interoperability due to class-heterogeneity – a result of inconsistent labeling schemes and partial annotations. Therefore, it is challenging to leverage these datasets in aggregate to train models with a complete representation of abnormalities that may occur within the thorax. In this work, we propose surgical aggregation, a federated learning framework for aggregating knowledge from class-heterogeneous datasets and learn a model that can simultaneously predict the presence of all disease labels present across the datasets. We evaluate our method using simulated and real-world class-heterogeneous datasets across both independent and identically distributed (iid) and non-iid settings. Our results show that surgical aggregation outperforms current methods, has better generalizability, and is a crucial first step towards tackling class-heterogeneity in federated learning to facilitate the development of clinically-useful models using previously non-interoperable chest x-ray datasets.
{IEEEkeywords}
Federated learning, class-heterogeneity, generalizability, client drift, chest x-ray
1 Introduction
Deep learning has revolutionized the field of radiology in recent years, especially for its utility in screening life-threatening conditions using chest x-rays (CXRs) [1]. As a result, numerous large-scale CXR datasets have been released for the purpose of training high-performing models with expert-level detection of diseases in the thorax [2, 3, 4, 5]. Since each dataset is curated at different sites using different acquisition parameters and image processing techniques, large-scale CXR datasets are heterogeneous and not independently and identically distributed (iid). Moreover, these datasets are class-heterogeneous due to inconsistent labeling schemes and partial annotations. In other words, each dataset focuses only on detecting a subset of findings that could be present in a patient and annotations present in one dataset may not be present in another – limiting their interoperability. For example, the NIH Chest X-ray 14 and CheXpert datasets contain 14 and 13 disease labels respectively, with only seven in common, making it challenging to leverage these datasets in aggregate to train large-scale models [5, 2].
Interoperability between these datasets is crucial as models trained on larger, diverse, multi-institutional datasets yield higher performance and better generalizability [6, 7, 8]. In the context of bias and fairness, models trained on multi-institutional datasets tend to the least biased towards under-served patient populations [9]. Therefore, there is a critically unmet need to address interoperability between large-scale CXR datasets, in order to train clinically-useful models with a complete representation of abnormalities that may occur in the thorax – even with the presence of class-heterogeneity.
Figure 1: An overview of surgical aggregation. Suppose two sites are curating CXR datasets to train models for their own niche applications; the first site works on cardiomegaly and pneumonia, while the second works on effusion and pneumonia. Using surgical aggregation, the central server aggregates ”knowledge” from each site for their respective tasks to learn a global model that can detect all three tasks simultaneously.
In this work, we propose surgical aggregation, a semi-supervised federated learning framework for aggregating knowledge from class-heterogeneous datasets (Fig. 1). In the aforementioned example, surgical aggregation can leverage the NIH and CheXpert datasets in aggregate to train a large-scale model that can simultaneously predict the presence of all 20 unique disease labels across both datasets – while performing as well as models trained individually on both datasets. We validate our framework by conducting a series of empirical experiments with varying levels of class- and data-heterogeneity to evaluate their stability using simulated iid datasets. Finally, we evaluate the efficacy of our framework in real-world settings using two large-scale non-iid datasets. Our main contributions are three-fold:
  1. 1.
    We propose a federated class-heterogeneous framework for training high-performing models that learn a complete representation of disease labels present across class-heterogeneous datasets.
  2. 2.
    Our extensive validation shows that surgical aggregation is stable to class- and data-heterogeneity and outperforms existing approaches.
  3. 3.
    We show that our framework has the potential to facilitate the development of large-scale models using previously non-interoperable datasets.
Our methods are detailed in Section 3. Convergence analysis for surgical aggregation is in Section 3.1.3. The experimental design is detailed in Section 4. Our results are detailed in Section 5 and discussed in Section 6. All our implementations are available on https://github.com/UM2ii/SurgicalAggregation.
2 Related Work
The prevalent approach to leverage knowledge from the myriad of CXR datasets is training models individually on each dataset. While they yield high-performance in isolation, individually trained models often do not generalize well in the real-world due to domain shift [10, 11, 12, 7, 8]. Cohen et al. [6] demonstrated that the poor generalizability of individually trained models stemmed from not just shift in imaging data, but also from shift in labels due to factors like inconsistent labeling schemes and clinical biases. Furthermore, this siloed approach would result in an ever increasing number of models as more datasets get curated in the future for novel tasks (e.g., detection of COVID-19).
Various groups have explored the idea of leveraging multiple datasets in aggregate to yield high-performing and highly-generalizable models [8, 7, 6, 9]. However, aggregating multi-site data in a centralized training setup is impractical due to patient data privacy. Federated learning (FL) was proposed as promising collaborative learning paradigm to mitigate this problem. Consisting of a central server with multiple clients, FL can train a global model on distributed datasets by aggregating ”knowledge” (i.e., model weights) learned by local models at each client using a strategy like FedAvg [13, 14]. It has enabled training of large-scale models using medical imaging data distributed across multiple institutions without sharing sensitive patient data [14, 15, 16]. However, class- and data-heterogeneity pose a significant challenge to FL.
Figure 2: Illustration of client drift in a FL setup in non-iid settings. As client model parameters move towards the client optima, the global model parameters drift away from the global optima for the network.
Due to data-heterogeneity, local models may learn features that differ significantly from each other (Fig. 2). Thus, different strategies have been proposed in literature for aggregating knowledge from non-iid datasets [17, 18, 19, 20, 21]. However, there is a significant gap in literature addressing class-heterogeneity. Most approaches focus on either detecting 1) the presence of an abnormality, or 2) the disease labels shared across all datasets [8, 7, 6, 9]. Furthermore, most work by other groups in leveraging class-heterogeneous datasets to train models has been in a centralized setup [22, 23, 24].
Kulkarni et al. [25] explored a personalized FL (PFL) approach using two class-heterogeneous CXR datasets. However, PFL requires each client to learn a unique model and potentially results in ambiguous predictions from client-to-client and poor generalizability in the real-world [26, 6]. Gong et al. [27] proposed an ensemble attention distillation-based FL approach to train a global model using NIH and CheXpert datasets. However, this approach learns a multi-class classifier, i.e., the model can predict the presence of only one disease – even if other diseases may be present. Furthermore, an additional dataset is required for distillation at the central server. Dong et al. [28] developed a class-heterogeneous FL framework that used local consistency regularization to generate pseudo-labels for non-local classes at a client. However, this approach is dependent on the representation learned by the unsupervised pre-training step to generate low-uncertainty pseudo-labels – a crucial consideration for non-iid settings.
For segmentation, Miao et al. [29] proposed a class-heterogeneous FL framework that corrects local optimization by providing direction of gradient for non-local tasks using a modified cross-entropy loss. While segmentation is a fundamentally different problem than classification, the authors similarly noted a lack of prior literature for federated class-heterogeneous learning frameworks that leverage multiple datasets in aggregate to train a singular multi-task model.
Therefore, our proposed framework is different from prior work in three major aspects. First, our framework is capable of training a singular multi-label global model that can simultaneously predict the presence of more than one disease. Second, our framework is semi-supervised and does not require an additional dataset. Finally and most importantly, our framework does not require clients to know the tasks nor share the same tasks of other clients to participate in the network – a core assumption of most FL techniques [14, 15, 16].
3 Methods
3.1 Surgical Aggregation
Surgical aggregation is a semi-supervised, model and task-agnostic FL framework for ”stitching”, i.e., selectively aggregating, task-specific weights from all clients to construct a global task block for the global model – even with the presence of class-heterogeneity in the network. Our framework is built upon the premise that our method can leverage partially annotated class-heterogeneous datasets to learn a sufficiently good approximation of a model trained with a hypothetical fully annotated equivalent.
3.1.1 Problem Setup
Let us consider a FL setup comprising of $K\in\mathbb{N}$ clients (indexed by $k$) training for $T\in\mathbb{N}$ epochs with communication after every $E\in\mathbb{N}$ epochs, resulting in $\lfloor T/E\rfloor$ communication rounds. Furthermore, each client $k$ has training data $\{\mathbf{x}_{k},y_{k}\}$ with $n_{k}$ samples.
Definition 1. We define a multi-label classifier as comprising of: 1) A representation block $F$ consisting of feature extraction layers, and 2) a task block $G$ consisting of fully-connected layers that interpret extracted features to predict classes, such that $\hat{y}=(G\circ F)(\mathbf{w};\mathbf{x})$ for data $\{\mathbf{x},y\}$.
Constraint 1. Let $C$ denote the complete set of classes distributed across the network and $C_{k}$ be the set of classes for client $k$, such that $C_{k}\subseteq C$ and $C=\bigcup_{k=1}^{K}C_{k}$.
Most FL approaches operate under the constraint that all classes are shared by all clients, i.e., $\forall c\in C_{i}$, $c\in C_{j}\ \forall i\neq j$. However, class-heterogeneity necessitates a broader definition:
Definition 2. We define class-heterogeneity as the case that covers: 1) Classes shared by all clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \forall i\neq j$, 2) Classes partially shared by some clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \exists i\neq j$, and 3) Classes unique to a client in the network, i.e., $\exists c\in C_{i}$ such that $c\not\in C_{j}\ \forall i\neq j$.
3.1.2 Algorithm Overview
Figure 3: Illustration of the surgical aggregation framework constructing a singular global task block using task-specific weights from each client in the network with task block aggregation strategy $g$ (e.g., mean).
At the start of each communication round, clients train locally for $E$ epochs before communicating the updated weights to the central server. The central server aggregates the weights of all layers in the representation block using strategy $f$ (e.g., FedAvg). Our framework uses a modified aggregation strategy to construct a singular task block using task-specific weights from each client in the network.
As shown in Fig. 3, if a class is shared across one or more clients, the weights corresponding to the class in the task block are aggregated using strategy $g$. In our implementation, we define $g$ as the mean of task-specific weights. However, if a class is not shared with any other client, the weights remain unchanged. This strategy enables the selective aggregation of weights applied to the input of task block to construct a global task block to predict all $C$ tasks in the network. The surgical aggregation framework is detailed in Algorithm 1.
Input: $K$ clients indexed by $k$ with local tasks $C_{k}$, total epochs $T$, epochs before communication $E$, model layers indexed by $l$, initialized model parameters $\mathbf{w}_{0,k}^{(l)}$, representation block strategy $f$, task block strategy $g$;
for each $\text{each epoch}\ t\in\{1,2,...,T\}$ do
       for each $\text{each client}\ k\ \text{and each layer}\ l$ do
             $\mathbf{w}_{t,k}^{(l)}\leftarrow SGD(\mathbf{w}_{t-1,k}^{(l)})$
       end for
[…]
             for each $\text{each client}\ k\ \text{and each layer}\ l$ do
[…]
                         $\mathbf{w}_{t,k}^{(l)}\leftarrow f(\mathbf{w}_{t}^{(l)})$
                   else
                         for each class $\text{each class c}\in C_{k}$ do
                               $(\mathbf{w}_{t,k}^{(l)})_{c}\leftarrow g((\mathbf{w}_{t}^{(l)})_{c})$
                         end for
                   end if
             end for
       end if
end for
Algorithm 1 Surgical Aggregation
3.1.3 Convergence Analysis
Let us consider a multi-label classifier with initialized parameters $\mathbf{w}_{0,k}$ updated using gradient descent. We primarily focus this convergence analysis on the modified strategy for task-specific weights in our method.
Definition 3. We define task-specific weights in the task block $G$ as $(\mathbf{w}_{t,k}^{(G)})_{c}$, i.e., the weights corresponding to epoch $t\in[T]$ and client $k$ for the class $c\in C_{k}$.
We analyze the convergence of surgical aggregation under scenarios of class-heterogeneity described in Definition 2.
Let $(G\circ F)(\mathbf{w})$ be the global model, while $(G\circ F)_{k}(\mathbf{w})$ be the local model for client $k$, such that our distributed optimization problem is defined as:
$\min_{\mathbf{w}}\left\{(G\circ F)^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}(G\circ F)_{k}^{*}(\mathbf{w})\right\}$ (1)
where the local objective $(G\circ F)_{k}^{*}$ is defined by:
$(G\circ F)_{k}^{*}\triangleq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$ (2)
where $L$ is the loss function (e.g., binary cross-entropy) [18].
Lemma 1. The optimization of the local objective $(G\circ F)^{*}_{k}$ for client $k$ is controlled by the optimization of $(G\circ F)^{*}_{k}$ on each individual class $c\in C_{k}$.
Proof: For multi-label classifiers, the loss at client $k$ is computed as the mean loss for each class $c\in C_{k}$ using loss function $L$, such that:
$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (3)
where $(y_{k,i})_{c}$ is the ground truth for class $c$ and sample $i\in[n_{k}]$ at client $k$. Therefore, to minimize $L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$, $\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ must be minimized for all $c\in C_{k}$. In other words, optimizing $(G\circ F)^{*}_{k}$ on each individual class $c\in C_{k}$ ensures optimization of the local objective $(G\circ F)^{*}_{k}$.
Furthermore, the global objective $(G\circ F)^{*}$ can be interpreted as the optimization of two concurrent sub-objectives:
  1. 1.
    Optimizing the global representation block $F$ by locally optimizing client representation blocks $F_{k}$, such that:
    $\min_{\mathbf{w}}\left\{F^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}^{*}(\mathbf{w})\right\}$ (4)
  2. 2.
    Optimizing the global task block $G$ by locally optimizing client task blocks $G_{k}$, such that:
    $\min_{\mathbf{w}}\left\{G^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}G_{k}^{*}(\mathbf{w})\right\}$ (5)
where the local sub-objectives $F_{k}^{*}=G_{k}^{*}\triangleq(G\circ F)_{k}^{*}$.
Lemma 2. The optimization of the global sub-objective $F^{*}$ is independent of class-heterogeneity present in the network.
Proof: Like conventional FL approaches, in our approach the representation block weights $\mathbf{w}_{k,t}^{(F)}$ are always shared with all clients at the end of each communication round, such that:
$\mathbf{w}_{t,k}^{(F)}=f(\mathbf{w}_{t}^{(F)})$ (6)
In this case, the convergence of $\mathbf{w}_{k,t}^{(F)}$ is well-established in prior literature [18, 17]. However, it is not clear whether class-heterogeneity affects the convergence of $\mathbf{w}_{k,t}^{(F)}$. The optimization of the global sub-objective $F^{*}$ is controlled by the optimization of the local objective $(G\circ F)^{*}_{k}$. Therefore, the optimization of $F^{*}$ is ensured if $F_{k}^{*}\triangleq(G\circ F)^{*}_{k}$ is optimized. While the optimization of $(G\circ F)^{*}_{k}$ is dependent on classes $C_{k}$, it is independent of the optimization of classes $c\in C\backslash C_{k}$ (Lemma 1). In other words, the optimization of $F^{*}$ is driven by the optimization of $F_{k}^{*}$, regardless of any class-heterogeneity.
Since the task block weights are selectively aggregated, we analyze the convergence of the global task block (i.e., $G^{*}$) with respect to each individual class $c\in C$.
Lemma 3. The optimization of the global objective $(G\circ F)^{*}$ with respect to a class $c\in C$, shared by a subset of clients $K^{\prime}$, is controlled by the optimization of $(G\circ F)_{k}^{*}$ for all clients $k\in K^{\prime}$ on the class $c$.
Proof: Let $K^{\prime}$ be the subset of clients in the network sharing class $c\in C$. In our approach, the task-specific weights $(\mathbf{w}_{k,t}^{(G)})_{c}$ for clients $k\in K^{\prime}$ are shared with the clients at the end of each communication round, such that:
$(\mathbf{w}_{k,t}^{(G)})_{c}=g((\mathbf{w}_{t}^{(G)})_{c})$ (7)
The global sub-objective $G^{*}$ with respect to the shared class $c$ is controlled by the local optimization of $G_{k}^{*}$ on class $c$ for each clients $k\in K^{\prime}$. Using Lemma 2, we conclude that with respect to shared class $c$, the optimization of global objective $(G\circ F)^{*}$ is ensured. In other words, regardless of whether all or some clients share the class $c$, the $(G\circ F)^{*}$ is ensured to be optimized with respect to the class $c$.
We further analyze a special case for Lemma 3 where $|K^{\prime}|=1$, i.e., there is a class $c\in C$ unique to a client $k$.
Lemma 4. The optimization of global objective $(G\circ F)^{*}$ with respect to a class $c\in C$, unique to client $k$, i.e., it is not shared by any other client in the network, is controlled by the optimization of $(G\circ F)_{k}^{*}$ on class $c$.
Proof: The global sub-objective $G^{*}$ with respect to the the class $c$ is by definition optimized by the local optimization of $G_{k}^{*}$ on class $c$ unique to client $k$. Using Lemma 2, we conclude that with respect to the class $c$, the optimization of the global objective $(G\circ F)^{*}$ is ensured.
Putting together Lemmas 3 and 4, we prove that the optimization of global objective $(G\circ F)^{*}$ with respect to class $c\in C$ is always ensured and is derived from previously established convergence analyses, regardless of class-heterogeneity in the network [18, 17]. Therefore, using Lemma 1, we can confer:
Theorem 1. The optimization of global objective $(G\circ F)^{*}$ is ensured if local objective $(G\circ F)_{k}^{*}$ is optimized for all classes $c\in C_{k}$ using representation block strategy $f$ and task block strategy $g$ and optimization is independent of any class-heterogeneity present in the network.
To conclude, Theorem 1 says that regardless of the class-heterogeneity present in the network, the convergence of surgical aggregation is guaranteed.
3.2 Comparison with Baselines
We compare our method to two baselines, i.e., models trained with well-established approaches in literature that represent an upper bound of the expected model performance, but do not yield a singular global model.
3.2.1 Individually Trained Baseline
The individually trained baseline refers to models trained individually on each dataset in a centralized setup using standard deep learning.
3.2.2 PFL Baseline
The PFL baseline encapsulates the growing popularity of personalized FL techniques in literature, where each client learns personalized local models in favor of generalized global models [30, 31, 17]. We implement the PFL baseline such that clients share a representation block, while learning a personalized task block for local classes.
3.3 Comparison with Existing Methods
We compare our method to three existing methods for training models by leveraging multiple datasets in aggregate. Unlike the baselines, these methods yield models with a singular task block.
3.3.1 Centralized
The centralized method refers to the well-established approach to achieve higher generalizability, where multiple datasets are concatenated locally to train a model in a centralized setup using standard deep learning [8, 7, 6, 9]. All missing classes in the aggregated dataset are treated as negatives, i.e., $(y_{i})_{c}=0$ for all $c\in C\backslash C_{k}$.
3.3.2 Vanilla FL
The vanilla FL method is the conventional approach for FL where datasets are distributed across clients in the network [14, 15, 16]. Similar to the centralized method, missing classes in each dataset are treated as negatives.
3.3.3 FL with Partial Loss
The FL with partial loss method is a modification of the vanilla FL setup, inspired by methods for training multi-label classifiers with missing labels to tackle class-heterogeneity using a modified binary cross-entropy loss [22, 23, 24]. Jin et al. [22] defines the partial loss function for a client $k$ with model parameters $\mathbf{w}$, local classes $C_{k}\subseteq C$, and local training data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples as follows:
$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (8)
3.4 Implementation
3.4.1 Centralized Setup
We use DenseNet121, pre-trained with ImageNet, to train multi-label classifiers using binary cross-entropy (BCE) loss with batch size of 64. Prior to training, task blocks are pre-trained using transfer learning for 15 epochs with learning rate of 5e-3, while keeping representation block weights frozen. Then, the model is fine-tuned for 150 epochs with learning rate of 5e-5.
3.4.2 Federated Setup
We simulate a FL setup using the above methods with $T=150$ epochs and $E=1$ epochs before communication, resulting in 150 communication rounds. Convergence of the global model is determined by the mean of local validation BCE loss for each client (Eq. 1). For each FL approach, we consider the three FL aggregation strategies:
  • FedAvg, the de-facto FL strategy [13]
  • FedBN, a variation of FedAvg with local batch normalization [17]. Since FedBN is designed primarily for PFL tasks, it is only included for the PFL baseline model.
  • FedBN+, a modification of FedBN with pre-trained batch normalization statistics [32].
3.4.3 Image Preprocessing
All CXRs are downsampled to $224\times 224$, normalized between 0 and 1, and scaled to ImageNet statistics. Random augmentations (rotation, flip, zoom, and contrast) were applied to input images during training.
3.4.4 Environment
All models are trained and evaluated using TensorFlow (version 2.8.1) and CUDA (version 12.0) on four NVIDIA RTX A6000 GPUs.
4 Experimental Setup
4.1 Datasets
We use three large-scale CXR datasets that are class- and data-heterogeneous, as shown in Fig. 4.
Figure 4: Illustration of class- and data-heterogeneity between the NIH, CheXpert, and MIMIC-LT datasets. (a) Barplot showing class-heterogeneity between the three datasets as the frequency of occurrence of disease labels. The seven shared classes between the NIH and CheXpert datasets are highlighted in bold. The MIMIC-LT dataset contains all 20 classes. (b) Latent space representation showing data-heterogeneity between all three datasets using t-SNE [33].
4.1.1 NIH Chest X-Ray 14
The NIH Chest X-Ray 14 dataset consists of 14 disease labels with $n=112,120$ frontal CXRs from 30,805 patients [5]. We randomly divide the dataset into training (70%, $n=78,075$), validation (10%, $n=11,079$), and testing (20%, $n=22,966$) splits with no patient leakage.
4.1.2 CheXpert
The CheXpert dataset consists of 13 disease labels (seven shared with the NIH dataset) with $n=224,316$ CXRs from 65,240 patients [2]. All lateral images are discarded to yield $n=191,027$ frontal CXRs. Uncertain labels are treated as negatives. We randomly divide the dataset into training (70%, $n=133,638$), validation (10%, $n=18,855$), and testing (20%, $n=38,534$) splits with no patient leakage.
4.1.3 MIMIC-CXR-JPG
The MIMIC-CXR-JPG dataset is used as our external test set to evaluate the generalizability to unseen data distributions [34, 3]. The MIMIC dataset consists of 13 disease labels (same as CheXpert) with $n=377,110$ CXRs from 65,379 patients. All lateral images are discarded to yield $n=243,324$ frontal CXRs. To thoroughly validate our method, we use the CXR-LT expanded set of 26 classes for the MIMIC dataset that provides ground truth for the 20 classes present across the NIH and CheXpert datasets as our fully annotated test set [35, 34].
Figure 5: Mean AUROC scores for iid simulation experiments across varying levels of class- and data-heterogeneity on the held-out NIH test set (left) and external MIMIC-LT dataset (right). (a) Effect of number of clients in the network as a proxy for data-heterogeneity. (b) Effect of number of shared classes across all clients in the network as a proxy for class-heterogeneity.
4.2 Experiments
4.2.1 Simulation Experiments
We evaluate the effect of class- and data-heterogeneity using simulated datasets created by sampling iid subsets of the NIH dataset for each client, with varying levels of class- and data-heterogeneity, such that there is no patient leakage and all 14 classes from the NIH dataset appear at least once in each simulated dataset (Constraint 1).
Since each simulated dataset is an incomplete snapshot of the NIH dataset, we aim to empirically quantify how well our method approximates the performance of the NIH baseline, trained on fully annotated data (Section 3). We evaluate our hypothesis across two sub-experiments:
First, we study the effect of data-heterogeneity across seven runs, each with an increasing number of clients $K\in\{2,3,4,5,6,8,10\}$ as a proxy for data-heterogeneity. As the number of clients increases, the sample size of the simulated dataset at each client decreases by factor of $1/K$, thus increasing statistical heterogeneity between clients. Second, we study the effect of class-heterogeneity across of seven runs, each with an increasing number of shared classes $\phi\in\{0,2,6,10,12,13,14\}$ as a proxy for class-heterogeneity. We keep the number of clients $K=4$ and underlying imaging data at each client constant.
All models are evaluated on the held-out NIH test set and the external MIMIC-LT dataset. For all FL approaches, FedBN+ is used as the aggregation strategy. The PFL baseline is not included due to its inability to train singular multi-task models.
4.2.2 Real-World Experiments
In addition to evaluating the stability of our framework across varying levels of class- and data-heterogeneity, we evaluate the efficacy of our framework in the real-world using two non-iid large-scale CXR datasets – NIH and CheXpert – to conduct two sub-experiments.
In the first sub-experiment, we consider the scenario without class-heterogeneity, an ideal scenario for existing FL setups, where only on the seven shared classes between both datasets are included as ground truth annotations (see Fig. 4). In the second experiment, we consider all classes to encapsulate a real-world class- and data-heterogeneous problem by training a model with a complete representation of all 20 classes across both datasets. All models are evaluated on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset.
Figure 6: Average mean AUROC scores for evaluating the effect of data-heterogeneity on the held-out NIH test set and external MIMIC-LT dataset using simulated iid datasets. Surgical aggregation is compared with existing methods for each test set using paired t-tests.
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
4.3 Metrics and Statistical Analyses
Model performance in a multi-label scenario is quantified by the mean of area under the receiver operating characteristic curve (AUROC) scores for each class in the test set. If a class in the test set is absent from the set of classes learned by a model, the mean AUROC score is not defined.
For the simulation experiments, performance between models is compared using paired t-tests between the mean AUROC scores to test for significance. For real-world experiments, performance between models is compared using paired t-tests across the granular class-level AUROC scores to test for significance. In all cases, statistical significance was defined as $p<0.05$ and the difference between paired samples is normal, indicated by $p\geq 0.05$ using the Shapiro-Wilk test.
5 Results
5.1 Simulation Experiments
Figure 7: Average mean AUROC scores for evaluating the effect of class-heterogeneity on the held-out NIH test set and external MIMIC-LT dataset using simulated iid datasets. Surgical aggregation is compared with existing methods for each test set using paired t-tests.
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
5.1.1 Effect of Data-Heterogeneity
Across varying levels of data-heterogeneity, we observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p<0.005$). While FL with partial loss performs comparably to our method, the difference in average mean AUROCs (margin $0.004$) is statistically significant ($p=0.004$). On the external MIMIC-LT dataset, our method generalizes better with average mean AUROC $0.69\pm 0.01$, compared to all existing methods (all $p<0.03$). Results are detailed in Figs. 5(a) and 6.
5.1.2 Effect of Class-Heterogeneity
Across varying levels of class-heterogeneity, with constant data-heterogeneity, we observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p<0.03$). Similarly, the minor differences in average mean AUROCs between FL with partial loss and surgical aggregation are statistically significant (margin $0.004$, $p=0.03$). On the external MIMIC-LT dataset, our method generalizes better than all existing methods with average mean AUROC $0.69\pm 0.01$. Results are detailed in Figs. 5(b) and 7.
Finally, across both sub-experiments, our results indicate that even with varying levels of class- and data-heterogeneity in the network, surgical aggregation provides the closest approximation to the performance of the NIH baseline model (mean AUROC $0.82$, mean margin $0.017$), followed by FL with partial loss as a close second (mean margin $0.021$).
Table 1: Mean AUROC scores on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset for real-world experiment without class-heterogeneity. All values reported as Mean $\pm$ SD. Best performing method is highlighted in bold. Surgical aggregation is compared with existing methods for each test set using paired t-tests.
NIH CheXpert MIMIC-LT
NIH Baseline $0.84\pm 0.05^{\text{ns}}$ $0.70\pm 0.07^{**}$ $0.73\pm 0.10^{**}$
CheXpert Baseline $0.81\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
PFL Baseline
 FedAvg $0.83\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
 FedBN $0.83\pm 0.05^{*}$ $0.77\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
 FedBN+ $0.83\pm 0.06^{**}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
Central $0.82\pm 0.06^{**}$ $0.77\pm 0.07^{\text{ns}}$ $0.76\pm 0.10^{\text{ns}}$
Vanilla FL
 FedAvg $0.81\pm 0.05^{***}$ $0.74\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
 FedBN+ $0.84\pm 0.05^{*}$ $0.75\pm 0.08^{*}$ $0.75\pm 0.09^{\text{ns}}$
FL + PartialLoss
 FedAvg $0.82\pm 0.05^{***}$ $0.75\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
 FedBN+ $0.83\pm 0.06^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
FL + SurgAgg
 FedAvg $0.84\pm 0.05^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
 FedBN+ $\mathbf{0.84}\pm 0.06$ $\mathbf{0.77}\pm 0.08$ $\mathbf{0.76}\pm 0.10$
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
5.2 Real-World Experiments
5.2.1 Data-Heterogeneous Scenario
We observe that the purely data-heterogeneous scenario results in similar mean AUROCs across the held-out NIH and CheXpert test sets and the external MIMIC-LT dataset. This implies that surgical aggregation is comparable to current FL approaches in the absence of class-heterogeneity (Table 1).
Specifically, on the NIH test set, surgical aggregation performs comparably to the NIH baseline with mean AUROC $0.84\pm 0.06$ ($p=0.22$), while outperforming other baselines (all $p<0.03$) and existing methods (all $p<0.02$).
On the CheXpert test set, surgical aggregation performs comparably to the CheXpert baseline model with mean AUROC $0.77\pm 0.08$ ($p=0.51$). Our method performs on par with the PFL baselines (all $p>0.05$) and existing methods, while outperforming the NIH baseline ($p=0.002$)
Finally, on the external MIMIC-LT dataset, we observe that surgical aggregation’s generalizability to unseen data is on par with baselines with mean AUROC $0.76\pm 0.10$ (all $p>0.05$), with the exception of the NIH baseline ($p=0.009$). Similarly, our method performs comparably to the existing methods.
Table 2: Mean AUROC scores on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset for real-world experiment with class-heterogeneity. All values reported as Mean $\pm$ SD. Best performing method is highlighted in bold. Surgical aggregation is compared with existing methods for each test set using paired t-tests.
NIH CheXpert MIMIC-LT
NIH Baseline $0.82\pm 0.06^{\text{ns}}$
CheXpert Baseline - $0.74\pm 0.07^{*}$
PFL Baseline
 FedAvg $0.79\pm 0.07^{***}$ $0.73\pm 0.07^{***}$
 FedBN $0.81\pm 0.06^{\text{ns}}$ $0.74\pm 0.06^{*}$
 FedBN+ $0.81\pm 0.06^{\text{ns}}$ $0.73\pm 0.07^{*}$
Central $0.71\pm 0.13^{**}$ $0.72\pm 0.07^{***}$ $0.65\pm 0.10^{***}$
Vanilla FL
 FedAvg $0.70\pm 0.10^{***}$ $0.68\pm 0.08^{***}$ $0.64\pm 0.09^{***}$
 FedBN+ $0.71\pm 0.12^{***}$ $0.68\pm 0.09^{***}$ $0.66\pm 0.10^{***}$
FL + PartialLoss
 FedAvg $0.77\pm 0.09^{***}$ $0.68\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
 FedBN+ $0.78\pm 0.07^{***}$ $0.71\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
FL + SurgAgg
 FedAvg $0.80\pm 0.06^{*}$ $0.73\pm 0.07^{**}$ $0.70\pm 0.09^{\text{ns}}$
 FedBN+ $\mathbf{0.81}\pm 0.07$ $\mathbf{0.76}\pm 0.08$ $\mathbf{0.71}\pm 0.10$
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
5.2.2 Class- and Data-Heterogeneous Scenario
Table 3: Sub-analysis of mean AUROC scores for shared and unique classes (Fig. 4) on the held-out NIH and CheXpert test sets for real-world experiment with class-heterogeneity. All values reported as Mean $\pm$ SD. Best performing method is highlighted in bold. Surgical aggregation is compared with existing methods for each test set and sub-group using paired t-tests.
NIH CheXpert
Shared Classes Unique Classes Shared Classes Unique Classes
NIH Baseline $0.84\pm 0.06^{\text{ns}}$ $0.80\pm 0.07^{\text{ns}}$ $0.70\pm 0.08^{**}$
CheXpert Baseline $0.81\pm 0.07^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.71\pm 0.05^{**}$
PFL Baseline $0.83\pm 0.06^{**}$ $0.80\pm 0.07^{\text{ns}}$ $0.76\pm 0.07^{\text{ns}}$ $0.70\pm 0.05^{**}$
Central $0.82\pm 0.05^{*}$ $0.60\pm 0.07^{***}$ $0.74\pm 0.07^{\text{ns}}$ $0.69\pm 0.06^{**}$
Vanilla FL $0.82\pm 0.06^{**}$ $0.60\pm 0.06^{***}$ $0.73\pm 0.09^{**}$ $0.62\pm 0.05^{**}$
FL + PartialLoss $0.82\pm 0.05^{*}$ $0.73\pm 0.06^{***}$ $0.74\pm 0.07^{*}$ $0.67\pm 0.04^{**}$
FL + SurgAgg $\mathbf{0.84}\pm 0.06$ $\mathbf{0.78}\pm 0.07$ $\mathbf{0.77}\pm 0.09$ $\mathbf{0.75}\pm 0.07$
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
Table 4: Sub-analysis of mean AUROC scores for NIH and CheXpert classes on the external MIMIC-LT dataset for real-world experiment with class-heterogeneity. All values reported as Mean $\pm$ SD. Best performing method is highlighted in bold. Surgical aggregation is compared with existing methods for each test set using paired t-tests.
MIMIC-LT
NIH Classes CheXpert Classes
NIH Baseline $**0.73\pm 0.10^{\text{**}}$
CheXpert Baseline - $0.67\pm 0.11^{*}$
PFL Baseline
 FedAvg $0.74\pm 0.11^{***}$ $0.69\pm 0.10^{\text{ns}}$
 FedBN $0.70\pm 0.09^{***}$ $0.67\pm 0.10^{\text{ns}}$
 FedBN+ $0.74\pm 0.10^{**}$ $0.68\pm 0.11^{\text{ns}}$
Central $0.74\pm 0.09^{**}$ $0.66\pm 0.11^{**}$
Vanilla FL
 FedAvg $0.72\pm 0.10^{***}$ $0.62\pm 0.06^{**}$
 FedBN+ $0.75\pm 0.10^{**}$ $0.63\pm 0.08^{*}$
FL + PartialLoss
 FedAvg $0.74\pm 0.09^{**}$ $0.66\pm 0.09^{*}$
 FedBN+ $0.74\pm 0.10^{**}$ $0.64\pm 0.10^{*}$
FL + SurgAgg
 FedAvg $0.75\pm 0.09^{\text{ns}}$ $0.68\pm 0.10^{\text{ns}}$
 FedBN+ $\mathbf{0.76}\pm 0.10$ $\mathbf{0.70}\pm 0.12$
ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
We observe that the introduction of class-heterogeneity results in a severe degradation of model performance for the existing methods, while the baseline methods demonstrate high-performance. Our results show that surgical aggregation continues to yield high-performance (Table 2).
Specifically, on the NIH test set, surgical aggregation performs comparably to the baseline models with mean AUROC $0.81\pm 0.07$ (all $p>0.05$), apart from the FedAvg PFL baseline ($p<0.001$). Furthermore, our method outperforms the existing methods (all $p<0.002$). On the CheXpert test set, surgical aggregation outperforms all baselines (all $p<0.03$) and existing methods (all $p<0.004$) with mean AUROC $0.76\pm 0.08$. Finally, on the external MIMIC-LT dataset, surgical aggregation outperforms all existing methods with mean AUROC $0.71\pm 0.10$ (all $p<0.001$).
We further explore the performance on a granular level by measuring the mean AUROC scores of the shared and the unique classes for the held-out NIH and CheXpert test sets (Table 3). We observe that the existing methods perform poorly on the unique classes when compared to the baselines. In stark contrast, surgical aggregation demonstrates high-performance on the unique classes. On the NIH test set, our method performs on par with the baselines ($p>0.05$), but on the CheXpert test set, surgical aggregation performs comparably to the baselines on the shared classes (both $p>0.05$), while outperforming them on the unique classes (both $p<0.008$).
Since the mean AUROC scores for baselines are not defined on the expanded MIMIC-LT classes, we explore generalizability by measuring mean AUROC scores of the NIH and CheXpert classes separately on the external MIMIC-LT dataset (Table 4). We observe that surgical aggregation generalizes better than the baselines and existing methods on NIH classes with mean AUROC $0.76\pm 0.10$ (all $p<0.006$). Similarly, on CheXpert classes, our method yields mean AUROC $0.70\pm 0.12$ and outperforms the existing methods (all $p<0.02$), while performing comparably to the baselines (all $p>0.05$), except the CheXpert baseline ($p=0.04$).
6 Discussion
Although the release of countless CXR datasets has spearheaded the utility of deep learning for expert-level detection of diseases in the thorax, the lack of interoperability due to class-heterogeneity has limited their utility beyond training ”narrowly” focused models. Therefore, there is a significant gap in literature for methods that address class-heterogeneity and enable training of large-scale models by leveraging non-interoperable datasets in aggregate.
Our results demonstrate that surgical aggregation enables interoperability between distributed class-heterogeneous datasets to collaboratively develop a global multi-task model. In the absence of class-heterogeneity, surgical aggregation performs on par with well-established approaches in literature. However, in the presence of class-heterogeneity, our method continues to maintain high-performance and performs on par with, or in some cases outperforms, the baselines. More importantly, in the presence of class-heterogeneity, surgical aggregation scales well across both iid and non-iid settings. On the other hand, existing methods fall short and scale poorly in non-iid settings. For example, while FL with partial loss is a close second in terms of performance with iid data, its performance significantly degrades with non-iid data. This is primarily due to feature shift between task-specific weights at each client as local updates from a client for non-local classes leads to client drift and degradation of performance on non-shared classes [17, 6]. Therefore, the selective aggregation of task-specific weights is critical to tackle class-heterogeneity in non-iid settings and is the crux of our method.
Another advantage of surgical aggregation is its ability to learn a singular multi-task global model that can predict the complete set of classes observed in the network. In stark contrast, the PFL paradigm entails $K$ clients learning $K$ personalized models. While it is popular approach to tackle non-iid data, PFL has interesting implications in the clinical environment. First, each personalized model will yield different predictions for the same CXR, thus, raising the question of which model’s prediction should be trusted. Second, PFL emphasizes local optimization over generalization, thus making personalized models susceptible to poor generalizability on unseen data and shift in labeling schemes [26, 6].
These advantages enable surgical aggregation to not just facilitate greater collaboration between institutes focusing on different tasks in the same domain, but also enable the countless ”narrowly” focused datasets (e.g., those curated for competitions) to be leveraged to collectively train clinically-useful models with a complete representation of abnormalities.
Despite our results demonstrating the advantages of surgical aggregation, our work has certain limitations. First, surgical aggregation is designed primarily for multi-label classifiers and may not scale for medical image segmentation – a fundamentally different and more difficult problem to solve. Second, due to limited literature for class-heterogeneous settings in FL, there are no established benchmarks to compare our method against beyond external validation. Third, our method requires all clients to share the same architecture for the representation block, an important consideration for clients with different computational resources. Finally, the empirical experiments are conducted on a simulated FL setup – lacking validation in real-world FL networks. For future work, we intend to explore solutions for these limitations and further validate our framework for real-world applications.
7 Conclusion
Surgical aggregation is a crucial first step towards tackling class-heterogeneity in FL, especially between large-scale CXR datasets. Our results show that surgical aggregation can train high-performing models that simultaneously predict the presence of all disease labels present across the datasets and facilitate the development of highly-generalizable models using previously non-interoperable CXR datasets.

### v6

@@
-Surgical Aggregation: Federated Class-Heterogeneous Learning
+From Isolation to Collaboration: Federated Class-Heterogeneous Learning for Chest X-Ray Classification
 Abstract
-The release of numerous chest x-ray datasets has spearheaded the development of deep learning models with expert-level performance.
-However, they have limited interoperability due to class-heterogeneity – a result of inconsistent labeling schemes and partial annotations.
-Therefore, it is challenging to leverage these datasets in aggregate to train models with a complete representation of abnormalities that may occur within the thorax.
-In this work, we propose surgical aggregation, a federated learning framework for aggregating knowledge from class-heterogeneous datasets and learn a model that can simultaneously predict the presence of all disease labels present across the datasets.
-We evaluate our method using simulated and real-world class-heterogeneous datasets across both independent and identically distributed (iid) and non-iid settings.
-Our results show that surgical aggregation outperforms current methods, has better generalizability, and is a crucial first step towards tackling class-heterogeneity in federated learning to facilitate the development of clinically-useful models using previously non-interoperable chest x-ray datasets.
-{IEEEkeywords}
-Federated learning, class-heterogeneity, generalizability, client drift, chest x-ray
+Federated learning (FL) is a promising paradigm to collaboratively train a global chest x-ray (CXR) classification model using distributed datasets while preserving patient privacy.
+A significant, yet relatively underexplored, challenge in FL is class-heterogeneity, where clients have different sets of classes.
+We propose surgical aggregation, a FL method that uses selective aggregation to collaboratively train a global model using distributed, class-heterogeneous datasets.
+Unlike other methods, our method does not rely on the assumption that clients share the same classes as other clients, know the classes of other clients, or have access to a fully annotated dataset.
+We evaluate surgical aggregation using class-heterogeneous CXR datasets across IID and non-IID settings.
+Our results show that our method outperforms current methods and has better generalizability.
+keywords:
+Federated learning, Multi-label classification, Class-heterogeneity, Chest X-ray
+Data and Code Availability
+All code and data used in this study is publicly available.
+The NIH Chest X-ray 14 dataset (https://nihcc.app.box.com/v/ChestXray-NIHCC) is open access.
+The CheXpert (https://stanfordmlgroup.github.io/competitions/chexpert/) and MIMIC-CXR-JPG (https://physionet.org/content/mimic-cxr-jpg/2.1.0/) datasets are credentialed access.
+Our code is available at https://github.com/BioIntelligence-Lab/SurgicalAggregation.
+Institutional Review Board (IRB)
+This retrospective study used publicly available datasets and was acknowledged by our IRB as exempt.
 1 Introduction
-Deep learning has revolutionized the field of radiology in recent years, especially for its utility in screening life-threatening conditions using chest x-rays (CXRs) [1].
-As a result, numerous large-scale CXR datasets have been released for the purpose of training high-performing models with expert-level detection of diseases in the thorax [2, 3, 4, 5].
-Since each dataset is curated at different sites using different acquisition parameters and image processing techniques, large-scale CXR datasets are heterogeneous and not independently and identically distributed (iid).
-Moreover, these datasets are class-heterogeneous due to inconsistent labeling schemes and partial annotations.
-In other words, each dataset focuses only on detecting a subset of findings that could be present in a patient and annotations present in one dataset may not be present in another – limiting their interoperability.
-For example, the NIH Chest X-ray 14 and CheXpert datasets contain 14 and 13 disease labels respectively, with only seven in common, making it challenging to leverage these datasets in aggregate to train large-scale models [5, 2].
-Interoperability between these datasets is crucial as models trained on larger, diverse, multi-institutional datasets yield higher performance and better generalizability [6, 7, 8].
-In the context of bias and fairness, models trained on multi-institutional datasets tend to the least biased towards under-served patient populations [9].
-Therefore, there is a critically unmet need to address interoperability between large-scale CXR datasets, in order to train clinically-useful models with a complete representation of abnormalities that may occur in the thorax – even with the presence of class-heterogeneity.
+Federated learning (FL) has emerged as a promising way to collaboratively train deep learning (DL) models for chest x-ray (CXR) classification using private datasets distributed across different institutions (Chowdhury et al., 2022; Sheller et al., 2020).
+Consisting of a central server and clients, FL can collaboratively train a global model by aggregating the weights of local models communicated by each client using a strategy (e.g., FedAvg) while preserving patient privacy (McMahan et al., 2017).
+A fundamental challenge in FL is heterogeneity, where differences in data distribution among clients in a network impacts model optimization (Li et al., 2020).
+This is relevant in medical imaging because each institution uses different acquisition parameters and image processing techniques, resulting in CXR datasets that are heterogeneous and not independently and identically distributed (IID).
+Moreover, each institution specializes in different tasks due to the costly and time-consuming process of manual annotation, resulting in partially annotated, class-heterogeneous datasets (Miao et al., 2023).
+In other words, annotations present in one dataset may not be present in another.
+For example, the NIH Chest X-Ray 14 and CheXpert datasets contain 14 and 13 disease labels respectively, with only seven in common (Wang et al., 2017; Irvin et al., 2019).
+Despite its prevalence in medical imaging, class-heterogeneity is a relatively underexplored topic in FL literature.
+While several groups have proposed techniques to tackle heterogeneity, they assume that all clients share the same set of classes (Li et al., 2021; Li and Wang, 2019; Karimireddy et al., 2020).
+However, in the real world, this may not always be the case.
+While some groups have attempted to solve class-heterogeneity, their techniques assume that clients know the classes of other clients (Miao et al., 2023; Dong et al., 2022) or use knowledge distillation with a public dataset in the central server that is either fully annotated (Gudur and Perepu, 2021) or unlabeled (Gong et al., 2021, 2022).
+Others avoid this challenge using personalized FL (PFL) techniques with client-specific classification heads (Kanhere et al., 2024; Kulkarni et al., 2022).
 Figure 1: An overview of surgical aggregation.
-Suppose two sites are curating CXR datasets to train models for their own niche applications; the first site works on cardiomegaly and pneumonia, while the second works on effusion and pneumonia.
-Using surgical aggregation, the central server aggregates ”knowledge” from each site for their respective tasks to learn a global model that can detect all three tasks simultaneously.
-In this work, we propose surgical aggregation, a semi-supervised federated learning framework for aggregating knowledge from class-heterogeneous datasets (Fig.
-1).
-In the aforementioned example, surgical aggregation can leverage the NIH and CheXpert datasets in aggregate to train a large-scale model that can simultaneously predict the presence of all 20 unique disease labels across both datasets – while performing as well as models trained individually on both datasets.
-We validate our framework by conducting a series of empirical experiments with varying levels of class- and data-heterogeneity to evaluate their stability using simulated iid datasets.
-Finally, we evaluate the efficacy of our framework in real-world settings using two large-scale non-iid datasets.
-Our main contributions are three-fold:
-  1. 1.
-    We propose a federated class-heterogeneous framework for training high-performing models that learn a complete representation of disease labels present across class-heterogeneous datasets.
-  2. 2.
-    Our extensive validation shows that surgical aggregation is stable to class- and data-heterogeneity and outperforms existing approaches.
-  3. 3.
-    We show that our framework has the potential to facilitate the development of large-scale models using previously non-interoperable datasets.
-Our methods are detailed in Section 3.
-Convergence analysis for surgical aggregation is in Section 3.1.3.
-The experimental design is detailed in Section 4.
-Our results are detailed in Section 5 and discussed in Section 6.
-All our implementations are available on https://github.com/UM2ii/SurgicalAggregation.
-2 Related Work
-The prevalent approach to leverage knowledge from the myriad of CXR datasets is training models individually on each dataset.
-While they yield high-performance in isolation, individually trained models often do not generalize well in the real-world due to domain shift [10, 11, 12, 7, 8].
-Cohen et al.
-[6] demonstrated that the poor generalizability of individually trained models stemmed from not just shift in imaging data, but also from shift in labels due to factors like inconsistent labeling schemes and clinical biases.
-Furthermore, this siloed approach would result in an ever increasing number of models as more datasets get curated in the future for novel tasks (e.g., detection of COVID-19).
-Various groups have explored the idea of leveraging multiple datasets in aggregate to yield high-performing and highly-generalizable models [8, 7, 6, 9].
-However, aggregating multi-site data in a centralized training setup is impractical due to patient data privacy.
-Federated learning (FL) was proposed as promising collaborative learning paradigm to mitigate this problem.
-Consisting of a central server with multiple clients, FL can train a global model on distributed datasets by aggregating ”knowledge” (i.e., model weights) learned by local models at each client using a strategy like FedAvg [13, 14].
-It has enabled training of large-scale models using medical imaging data distributed across multiple institutions without sharing sensitive patient data [14, 15, 16].
-However, class- and data-heterogeneity pose a significant challenge to FL.
-Figure 2: Illustration of client drift in a FL setup in non-iid settings.
-As client model parameters move towards the client optima, the global model parameters drift away from the global optima for the network.
-Due to data-heterogeneity, local models may learn features that differ significantly from each other (Fig.
-2).
-Thus, different strategies have been proposed in literature for aggregating knowledge from non-iid datasets [17, 18, 19, 20, 21].
-However, there is a significant gap in literature addressing class-heterogeneity.
-Most approaches focus on either detecting 1) the presence of an abnormality, or 2) the disease labels shared across all datasets [8, 7, 6, 9].
-Furthermore, most work by other groups in leveraging class-heterogeneous datasets to train models has been in a centralized setup [22, 23, 24].
-Kulkarni et al.
-[25] explored a personalized FL (PFL) approach using two class-heterogeneous CXR datasets.
-However, PFL requires each client to learn a unique model and potentially results in ambiguous predictions from client-to-client and poor generalizability in the real-world [26, 6].
-Gong et al.
-[27] proposed an ensemble attention distillation-based FL approach to train a global model using NIH and CheXpert datasets.
-However, this approach learns a multi-class classifier, i.e., the model can predict the presence of only one disease – even if other diseases may be present.
-Furthermore, an additional dataset is required for distillation at the central server.
-Dong et al.
-[28] developed a class-heterogeneous FL framework that used local consistency regularization to generate pseudo-labels for non-local classes at a client.
-However, this approach is dependent on the representation learned by the unsupervised pre-training step to generate low-uncertainty pseudo-labels – a crucial consideration for non-iid settings.
-For segmentation, Miao et al.
-[29] proposed a class-heterogeneous FL framework that corrects local optimization by providing direction of gradient for non-local tasks using a modified cross-entropy loss.
-While segmentation is a fundamentally different problem than classification, the authors similarly noted a lack of prior literature for federated class-heterogeneous learning frameworks that leverage multiple datasets in aggregate to train a singular multi-task model.
-Therefore, our proposed framework is different from prior work in three major aspects.
-First, our framework is capable of training a singular multi-label global model that can simultaneously predict the presence of more than one disease.
-Second, our framework is semi-supervised and does not require an additional dataset.
-Finally and most importantly, our framework does not require clients to know the tasks nor share the same tasks of other clients to participate in the network – a core assumption of most FL techniques [14, 15, 16].
+Suppose two sites want to collaboratively train a CXR classification model using FL.
+The first site works on cardiomegaly and pneumonia, while the second works on effusion and pneumonia.
+Using surgical aggregation, the central server can learn a global model to predict all three labels.
+[…]
 2 Methods
 2.1 Surgical Aggregation
-Surgical aggregation is a semi-supervised, model and task-agnostic FL framework for ”stitching”, i.e., selectively aggregating, task-specific weights from all clients to construct a global task block for the global model – even with the presence of class-heterogeneity in the network.
-Our framework is built upon the premise that our method can leverage partially annotated class-heterogeneous datasets to learn a sufficiently good approximation of a model trained with a hypothetical fully annotated equivalent.
+Surgical aggregation is based on the selective aggregation of weights communicated by clients to dynamically build a global classification head that predicts all classes available in the network from distributed, class-heterogeneous datasets.
+Our method does not rely on clients sharing the same set of classes or handling other non-local classes to cover three types of partial annotations: 1) Classes shared by all clients, 2) Classes shared by some clients, and 3) Classes unique to a client.
 2.1.1 Problem Setup
 Let us consider a FL setup comprising of $K\in\mathbb{N}$ clients (indexed by $k$) training for $T\in\mathbb{N}$ epochs with communication after every $E\in\mathbb{N}$ epochs, resulting in $\lfloor T/E\rfloor$ communication rounds.
-Furthermore, each client $k$ has training data $\{\mathbf{x}_{k},y_{k}\}$ with $n_{k}$ samples.
-Definition 1.
-We define a multi-label classifier as comprising of: 1) A representation block $F$ consisting of feature extraction layers, and 2) a task block $G$ consisting of fully-connected layers that interpret extracted features to predict classes, such that $\hat{y}=(G\circ F)(\mathbf{w};\mathbf{x})$ for data $\{\mathbf{x},y\}$.
-Constraint 1. Let $C$ denote the complete set of classes distributed across the network and $C_{k}$ be the set of classes for client $k$, such that $C_{k}\subseteq C$ and $C=\bigcup_{k=1}^{K}C_{k}$.
-Most FL approaches operate under the constraint that all classes are shared by all clients, i.e., $\forall c\in C_{i}$, $c\in C_{j}\ \forall i\neq j$.
-However, class-heterogeneity necessitates a broader definition:
-Definition 2.
-We define class-heterogeneity as the case that covers: 1) Classes shared by all clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \forall i\neq j$, 2) Classes partially shared by some clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \exists i\neq j$, and 3) Classes unique to a client in the network, i.e., $\exists c\in C_{i}$ such that $c\not\in C_{j}\ \forall i\neq j$.
-3.1.2 Algorithm Overview
-Figure 3: Illustration of the surgical aggregation framework constructing a singular global task block using task-specific weights from each client in the network with task block aggregation strategy $g$ (e.g., mean).
-At the start of each communication round, clients train locally for $E$ epochs before communicating the updated weights to the central server.
-The central server aggregates the weights of all layers in the representation block using strategy $f$ (e.g., FedAvg).
-Our framework uses a modified aggregation strategy to construct a singular task block using task-specific weights from each client in the network.
-As shown in Fig.
-3, if a class is shared across one or more clients, the weights corresponding to the class in the task block are aggregated using strategy $g$.
-In our implementation, we define $g$ as the mean of task-specific weights.
-However, if a class is not shared with any other client, the weights remain unchanged.
-This strategy enables the selective aggregation of weights applied to the input of task block to construct a global task block to predict all $C$ tasks in the network.
-The surgical aggregation framework is detailed in Algorithm 1.
-Input: $K$ clients indexed by $k$ with local tasks $C_{k}$, total epochs $T$, epochs before communication $E$, model layers indexed by $l$, initialized model parameters $\mathbf{w}_{0,k}^{(l)}$, representation block strategy $f$, task block strategy $g$;
+Let $C_{k}$ be the set of local classes for client $k$.
+We define $C$ as the complete set of classes distributed across the network, such that $C=\bigcup_{k=1}^{K}C_{k}$.
+Figure 2: Illustration of surgical aggregation dynamically building a global classification head using class-specific weights.
+We define a multi-label classifier $(G\circ F)$ as comprising of: 1) feature extractor $F$, and 2) classification head $G$.
+Here, $F$ can be any model backbone (e.g., DenseNet) and $G$ is a fully-connected layer.
+Suppose each client $k$ has local model $(G\circ F)_{k}$, where $F_{k}$ is the local feature extractor and $G_{k}$ is the local classification head.
+At an epoch $t$, we denote the weights of the local feature extractor as $\mathbf{w}_{t,k}^{(F)}$ and global feature extractor as $\mathbf{w}_{t}^{(F)}$.
+Similarly, we denote the weights of the local classification head as $\mathbf{w}_{t,k}^{(G)}\in\mathbb{R}^{N\times M_{k}}$ and global classification head as $\mathbf{w}_{t}^{(G)}\in\mathbb{R}^{N\times M}$.
+Here, $N$ is the number of input neurons to the fully-connected layer, $M$ and $M_{k}$ are the cardinality of $C$ and $C_{k}$ respectively (such that $M\geq M_{k}\ \forall k$).
+We define the class-specific weights of class $c\in C_{k}$ for client $k$ as column vector of $c$ in $\mathbf{w}_{t,k}^{(G)}$ and is denoted $(\mathbf{w}_{t,k}^{(G)})_{c}\in\mathbb{R}^{N}$.
+If $c\not\in C_{k}$, then $(\mathbf{w}_{t,k}^{(G)})_{c}=\mathbf{0}$.
+2.1.2 Overview of Surgical Aggregation
+At the start of each communication round, clients train locally using gradient descent for $E$ epochs on local classes before communicating the updated weights to the central server.
+To update the global model, we use a modified server update strategy comprising of feature extractor update and classification head update steps.
+Feature Extractor Update: We aggregate the weights of local feature extractors $\mathbf{w}_{t,k}^{(F)}$ using model aggregation strategy $f$.
+Here, $f$ can be any model aggregation strategy (e.g., FedAvg).
+For example, if $f$ is FedAvg, then the weights of the global feature extractor are defined as (McMahan et al., 2017):
+$\mathbf{w}_{t}^{(F)}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t,k}^{(l)}$ (1)
+Figure 3: Detailed illustration of surgical aggregation building the weights of the global classification head.
+[…]
+For class $c\in C$ present in $K_{c}\in\mathbb{N}$ clients (such that $K_{c}\leq K$), the global class-specific weights are computed by the mean of all local class-specific weights:
+$(\mathbf{w}_{t}^{(G)})_{c}=\frac{1}{K_{c}}\sum_{k=1}^{K}(\mathbf{w}_{t,k}^{(G)})_{c}$ (2)
+If the class $c$ is unique to a client (i.e. $K_{c}=1$), Equation 2 says that the global class-specific weights are equal to the local class-specific weights:
+$(\mathbf{w}_{t}^{(G)})_{c}=(\mathbf{w}_{t,k}^{(G)})_{c}$ (3)
+[…]
+At the end of the communication round, the aggregated $\mathbf{w}_{t}^{(F)}$ is communicated back to the clients.
+However, our method does not directly communicate the aggregated $\mathbf{w}_{t}^{(G)}$ back to the clients because it contains class-specific weights for all classes, including non-local classes $c\not\in C_{k}$ for a client $k$.
+Instead, our method reconstructs and communicates $\mathbf{w}_{t,k}^{(G)}$ to each clients $k$, using the updated $(\mathbf{w}_{t}^{(G)})_{c}$ for all $c\in C_{k}$.
+Input: $K$ clients indexed by $k$ with local classes $C_{k}$, total epochs $T$, epochs before communication $E$, initialized model parameters $\mathbf{w}_{0,k}$, model aggregation strategy $f$
 for each $\text{each epoch}\ t\in\{1,2,...,T\}$ do
-       for each $\text{each client}\ k\ \text{and each layer}\ l$ do
-             $\mathbf{w}_{t,k}^{(l)}\leftarrow SGD(\mathbf{w}_{t-1,k}^{(l)})$
+       // client update
+       for each $\text{each client}\ k$ do
+             $\mathbf{w}_{t,k}\leftarrow SGD(\mathbf{w}_{t-1,k})$
        end for
-[…]
-             for each $\text{each client}\ k\ \text{and each layer}\ l$ do
-[…]
-                         $\mathbf{w}_{t,k}^{(l)}\leftarrow f(\mathbf{w}_{t}^{(l)})$
-                   else
-                         for each class $\text{each class c}\in C_{k}$ do
-                               $(\mathbf{w}_{t,k}^{(l)})_{c}\leftarrow g((\mathbf{w}_{t}^{(l)})_{c})$
+      // server update
+       if $\text{mod}(t,E)=0$ then
+             // feature extractor update
+             $\mathbf{w}_{t}^{(F)}\leftarrow f(\{\mathbf{w}_{t,k}^{(F)}\ \text{: }\forall k\})$
+            // classification head update
+             for each $\text{each class}\ c\in C$ do
+                   $(\mathbf{w}_{t}^{(G)})_{c}\leftarrow mean(\{(\mathbf{w}_{t,k}^{(G)})_{c}\ \text{if }c\in C_{k}\ \text{: }\forall k\})$
              end for
-                   end if
+            // communicate to clients
+             for each $\text{each client}\ k$ do
+                   $\mathbf{w}_{t,k}^{(F)}\leftarrow\mathbf{w}_{t}^{(F)}$
+                   for each $\text{each class}\ c\in C_{k}$ do
+                         $(\mathbf{w}_{t,k}^{(G)})_{c}\leftarrow(\mathbf{w}_{t}^{(G)})_{c}$
+                   end for
              end for
@@
 Algorithm 1 Surgical Aggregation
-3.1.3 Convergence Analysis
-Let us consider a multi-label classifier with initialized parameters $\mathbf{w}_{0,k}$ updated using gradient descent.
-We primarily focus this convergence analysis on the modified strategy for task-specific weights in our method.
-Definition 3. We define task-specific weights in the task block $G$ as $(\mathbf{w}_{t,k}^{(G)})_{c}$, i.e., the weights corresponding to epoch $t\in[T]$ and client $k$ for the class $c\in C_{k}$.
-We analyze the convergence of surgical aggregation under scenarios of class-heterogeneity described in Definition 2.
-Let $(G\circ F)(\mathbf{w})$ be the global model, while $(G\circ F)_{k}(\mathbf{w})$ be the local model for client $k$, such that our distributed optimization problem is defined as:
-$\min_{\mathbf{w}}\left\{(G\circ F)^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}(G\circ F)_{k}^{*}(\mathbf{w})\right\}$ (1)
+[…]
+2.2 Datasets
+[…]
+Figure 4: Freqeuence of disease labels between NIH, CheXpert, and MIMIC datasets. The seven shared classes between the NIH and CheXpert datasets are bolded.
+NIH Chest X-Ray 14: It consists of 14 disease labels with $n=112,120$ frontal CXRs from 30,805 patients (Wang et al., 2017).
+We randomly divide the dataset into training (70%, $n=78,075$), validation (10%, $n=11,079$), and testing (20%, $n=22,966$) splits with no patient leakage.
+CheXpert: It consists of 13 disease labels (seven shared with NIH) with $n=224,316$ CXRs from 65,240 patients (Irvin et al., 2019).
+All lateral images are discarded to yield $n=191,027$ frontal CXRs.
+Uncertain labels are treated as negatives.
+We randomly divide the dataset into training (70%, $n=133,638$), validation (10%, $n=18,855$), and testing (20%, $n=38,534$) splits with no patient leakage.
+MIMIC-CXR-JPG: It consists of $n=377,110$ CXRs from 65,379 patients and is used as our external, fully annotated test set to evaluate generalizability (Goldberger et al., 2000; Johnson et al., 2019).
+All lateral images are discarded to yield $n=243,324$ frontal CXRs.
+The CXR-LT expanded set of 26 labels provides ground-truth for all 20 classes present in NIH and CheXpert datasets (Holste et al., 2023).
+Figure 5: Latent space representation of NIH, CheXpert, and MIMIC datasets using t-SNE (Van der Maaten and Hinton, 2008).
+2.3 Baselines
+We compare our method to two baselines that represent an upper bound of expected model performance, but do not yield a singular global model.
+Individually Trained: Models trained individually on each dataset in a centralized setup using standard DL.
+Personalized FL: PFL is a popular approach to FL, where each client learns a personalized local model in favor of generalized global models (Liang et al., 2020; Tan et al., 2022; Li et al., 2021).
+While clients share a feature extractor, they learn a personalized classification head for local classes.
+We also compare our method to three existing methods for training models with class-heterogeneous datasets that result in a single global model.
+Centralized: Multiple datasets are concatenated locally to train a model in a centralized setup using standard DL to achieve better generalizability (Cohen et al., 2020; Seyyed-Kalantari et al., 2021).
+All missing classes in the combined dataset are treated as negatives.
+For a client $k$ with local classes $C_{k}$, and data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$), $(y_{k,i})_{c}=0$ for all $c\not\in C_{k}$.
+Vanilla FL: Conventional approach for FL where datasets are distributed across clients in the network (Chowdhury et al., 2022; Sheller et al., 2020).
+Missing classes in each dataset are treated as negatives.
+FL with Partial Loss: A modification of the vanilla FL setup using a partial loss function that drops non-local classes when computing local losses (Jin and Ghahramani, 2002).
+For client $k$ with model parameters $\mathbf{w}_{t,k}$, local classes $C_{k}$, and private data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$), it is defined as:
+$L(\mathbf{w}_{t,k},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w}_{t,k},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (4)
+2.4 Experiments
+2.4.1 Ablation Study
+In this experiment, we evaluated the effect of 1) number of clients and 2) number of shared classes using simulated datasets.
+We sampled IID subsets of the NIH dataset to create simulated datasets for each client, such that there is no patient leakage and all 14 classes from the NIH dataset appear at least once in the network.
+Since each simulated dataset is an ”incomplete” snapshot of the NIH dataset, we evaluated the performance of surgical aggregation in relation to the NIH baseline model.
+Effect of number of clients: We evaluated the effect of number of clients across seven iterations, each with an increasing number of clients $K\in\{2,3,4,5,6,8,10\}$.
+We randomly selected values of $K$, such that as $K$ increases, the sample size of the simulated dataset at each client $k$ decreases by factor of $1/K$.
+Effect of number of shared classes: We evaluated the effect of number of shared classes as a proxy for class-heterogeneity in the network.
+Since the NIH dataset contains 14 classes, we randomly sampled seven possible iterations out of 14, each with an increasing number of shared classes $\{0,1,2,4,8,12,14\}$.
+The number of clients $K=4$ and underlying data at each client was constant.
+All models were evaluated on the held-out NIH test set and the external MIMIC dataset.
+We used FedBN+ for all FL approaches (Kulkarni et al., 2023).
+The PFL baseline was not included because it does not train a global model.
+2.4.2 Non-IID Experiments
+We evaluated surgical aggregation using the NIH and CheXpert datasets across two sub-experiments:
+[…]
+Class-Heterogeneous: We considered a non-IID and class-heterogeneous setting, where all 20 classes from both datasets were included as ground-truth annotations.
+All models were evaluated on the NIH and CheXpert test sets using local classes (14 and 13 respectively) and the external MIMIC dataset using all 20 classes.
+We also measured the disparity in performance between the shared and unique classes for the NIH and CheXpert datasets using their respective test set.
+Furthermore, we measured the performance on the NIH and CheXpert classes when evaluating with the external MIMIC test set.
+This enabled us to compare our method with the individually trained and PFL baselines.
+2.5 Implementation
+Centralized Setup: We trained ImageNet-pretrained DenseNet121 models using binary cross-entropy (BCE) loss and a batch size of 64.
+Prior to training, the classification heads were ”warmed up” using transfer learning for 15 epochs with a learning rate of 5e-3, while keeping feature extractor weights frozen.
+Then, the model was trained with all weights unfrozen for 150 epochs with a learning rate of 5e-5.
+Federated Setup: We simulated a FL setup with $T=150$ epochs and $E=1$ epochs before communication.
+Each client trained an ImageNet-pretrained DenseNet121 model using the methods described above.
+The classification heads of local models were similarly ”warmed up” prior to training.
+Global convergence was determined by the mean of local validation BCE loss for each client.
+We considered the three aggregation strategies: 1) FedAvg, the de-facto FL strategy (McMahan et al., 2017).
+2) FedBN, a variation of FedAvg with local batch normalization (Li et al., 2021).
+As a PFL strategy, FedBN is only included for the PFL baseline model.
+3) FedBN+, a modification of FedBN with pre-trained batch normalization statistics for global models (Kulkarni et al., 2023).
+In order to primarily focus on class-heterogeneity, we used the same hyperparameters for all models trained.
+All CXRs were downsampled to $224\times 224$, normalized between 0 and 1, and scaled to ImageNet statistics.
+Random augmentations (rotation, flip, zoom, and contrast) were applied to input images during training.
+All models were trained and evaluated using TensorFlow (version 2.8.1) and CUDA (version 12.0) on four NVIDIA RTX A6000 GPUs.
+Figure 6: Mean AUROC scores for ablation study evaluating (a) effect of number of clients, (b) effect of number of shared classes on the NIH (left) and MIMIC (right) test sets.
+Figure 7: Average mean AUROC scores for the effect of number of clients on NIH and MIMIC test sets. (ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+Figure 8: Average mean AUROC scores for the effect of number of shared classes on NIH and MIMIC test sets. (ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+2.6 Metrics and Statistical Analyses
+Model performance in a multi-label scenario is quantified by the mean of area under the receiver operating characteristic curve (AUROC) scores for each class in the test set.
+If a class in the test set is absent from the set of classes learned by a model, the mean AUROC score is not defined.
+All values are reported as Mean $\pm$ SD.
+For the ablation study, performance between models is compared using paired t-tests between the mean AUROC scores to test for significance.
+For real-world experiments, performance between models is compared using paired t-tests across the granular class-level AUROC scores to test for significance.
+In all cases, statistical significance was defined as $p<0.05$.
+3 Results
+3.1 Ablation Study
+Effect of number of clients: We observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p\leq 0.005$) (Figures 6(a) and 7).
+While FL with partial loss performs comparably to our method, the difference in average mean AUROCs ($\Delta=0.004$) is statistically significant ($p=0.004$).
+On the external MIMIC dataset, our method generalizes better with average mean AUROC $0.69\pm 0.01$, compared to all existing methods (all $p\leq 0.03$).
+Effect of number of shared clients: We observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p\leq 0.03$) (Figures 6(b) and 8).
+Similarly, the minor differences in average mean AUROCs between FL with partial loss and surgical aggregation are statistically significant ($\Delta=0.004$, $p=0.03$).
+On the external MIMIC dataset, our method generalizes better than all existing methods with average mean AUROC $0.69\pm 0.01$.
+Finally, our results indicate that even with varying levels of class- and data-heterogeneity in the network, surgical aggregation provides the closest approximation to the performance of the NIH baseline model (mean AUROC $0.82$, mean $\Delta=0.017$), followed by FL with partial loss (mean $\Delta=0.021$).
+Table 1: Mean AUROC scores on NIH, CheXpert, and MIMIC dataset for non-IID, non-class-heterogeneous experiment.
+Our method is compared with others using paired t-tests.
+Best performing method is bolded.
+NIH CheXpert MIMIC
+NIH $0.84\pm 0.05^{\text{ns}}$ $0.70\pm 0.07^{**}$ $0.73\pm 0.10^{**}$
+CheXpert $0.81\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
+PFL
+FedAvg $0.83\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
+FedBN $0.83\pm 0.05^{*}$ $0.77\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
+FedBN+ $0.83\pm 0.06^{**}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
+Central $0.82\pm 0.06^{**}$ $0.77\pm 0.07^{\text{ns}}$ $0.76\pm 0.10^{\text{ns}}$
+FL
+FedAvg $0.81\pm 0.05^{***}$ $0.74\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
+FedBN+ $0.84\pm 0.05^{*}$ $0.75\pm 0.08^{*}$ $0.75\pm 0.09^{\text{ns}}$
+FL + PL
+FedAvg $0.82\pm 0.05^{***}$ $0.75\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
+FedBN+ $0.83\pm 0.06^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
+FL + SA
+FedAvg $0.84\pm 0.05^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
+FedBN+ $\mathbf{0.84}\pm 0.06$ $\mathbf{0.77}\pm 0.08$ $\mathbf{0.76}\pm 0.10$
+PL: Partial Loss, SA: Surgical Aggregation
+(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+3.2 Non-IID Experiments
+[…]
+On the NIH test set, our method performs comparably to the NIH baseline with mean AUROC $0.84\pm 0.06$ ($p=0.22$), while outperforming other baselines (all $p\leq 0.03$) and existing methods (all $p\leq 0.02$).
+On the CheXpert test set, our method performs comparably to the CheXpert baseline model with mean AUROC $0.77\pm 0.08$ ($p=0.51$).
+Our method performs on par with the PFL baselines (all $p>0.05$) and existing methods, while outperforming the NIH baseline ($p=0.002$)
+Finally, on the external MIMIC dataset, we observe that surgical aggregation’s generalizability to unseen data is on par with baselines with mean AUROC $0.76\pm 0.10$ (all $p>0.05$), while outperforming the NIH baseline ($p=0.009$).
+Similarly, our method performs comparably to the existing methods.
+Table 2: Mean AUROC scores on NIH, CheXpert and MIMIC test sets for non-IID, class-heterogeneous experiment. Our method is compared with others using paired t-tests. Best performing method is bolded.
+NIH CheXpert MIMIC
+NIH $0.82\pm 0.06^{\text{ns}}$
+CheXpert - $0.74\pm 0.07^{*}$
+PFL
+FedAvg $0.79\pm 0.07^{***}$ $0.73\pm 0.07^{***}$
+FedBN $0.81\pm 0.06^{\text{ns}}$ $0.74\pm 0.06^{*}$
+FedBN+ $0.81\pm 0.06^{\text{ns}}$ $0.73\pm 0.07^{*}$
+Central $0.71\pm 0.13^{**}$ $0.72\pm 0.07^{***}$ $0.65\pm 0.10^{***}$
+FL
+FedAvg $0.70\pm 0.10^{***}$ $0.68\pm 0.08^{***}$ $0.64\pm 0.09^{***}$
+FedBN+ $0.71\pm 0.12^{***}$ $0.68\pm 0.09^{***}$ $0.66\pm 0.10^{***}$
+FL + PL
+FedAvg $0.77\pm 0.09^{***}$ $0.68\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
+FedBN+ $0.78\pm 0.07^{***}$ $0.71\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
+FL + SA
+FedAvg $0.80\pm 0.06^{*}$ $0.73\pm 0.07^{**}$ $0.70\pm 0.09^{\text{ns}}$
+FedBN+ $\mathbf{0.81}\pm 0.07$ $\mathbf{0.76}\pm 0.08$ $\mathbf{0.71}\pm 0.10$
+PL: Partial Loss, SA: Surgical Aggregation
+(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+[…]
+Table 3: Sub-analysis of mean AUROC scores for shared and unique classes of the NIH and CheXpert datasets for non-IID, class-heterogeneous experiment.
+Our method is compared with others using paired t-tests.
+Best performing method is bolded.
+NIH CheXpert
+Shared Classes Unique Classes Shared Classes Unique Classes
+NIH $0.84\pm 0.06^{\text{ns}}$ $0.80\pm 0.07^{\text{ns}}$ $0.70\pm 0.08^{**}$
+CheXpert $0.81\pm 0.07^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.71\pm 0.05^{**}$
+PFL $0.83\pm 0.06^{**}$ $0.80\pm 0.07^{\text{ns}}$ $0.76\pm 0.07^{\text{ns}}$ $0.70\pm 0.05^{**}$
+Central $0.82\pm 0.05^{*}$ $0.60\pm 0.07^{***}$ $0.74\pm 0.07^{\text{ns}}$ $0.69\pm 0.06^{**}$
+FL $0.82\pm 0.06^{**}$ $0.60\pm 0.06^{***}$ $0.73\pm 0.09^{**}$ $0.62\pm 0.05^{**}$
+FL + PL $0.82\pm 0.05^{*}$ $0.73\pm 0.06^{***}$ $0.74\pm 0.07^{*}$ $0.67\pm 0.04^{**}$
+FL + SA $\mathbf{0.84}\pm 0.06$ $\mathbf{0.78}\pm 0.07$ $\mathbf{0.77}\pm 0.09$ $\mathbf{0.75}\pm 0.07$
+PL: Partial Loss, SA: Surgical Aggregation
+(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+Table 4: Sub-analysis of mean AUROC scores for NIH and CheXpert classes on the MIMIC test for non-IID, class-heterogeneous experiment.
+Our method is compared with others using paired t-tests.
+Best performing method is bolded.
+MIMIC
+NIH Classes CheXpert Classes
+NIH $**0.73\pm 0.10^{\text{**}}$
+CheXpert - $0.67\pm 0.11^{*}$
+PFL
+FedAvg $0.74\pm 0.11^{***}$ $0.69\pm 0.10^{\text{ns}}$
+FedBN $0.70\pm 0.09^{***}$ $0.67\pm 0.10^{\text{ns}}$
+FedBN+ $0.74\pm 0.10^{**}$ $0.68\pm 0.11^{\text{ns}}$
+Central $0.74\pm 0.09^{**}$ $0.66\pm 0.11^{**}$
+FL
+FedAvg $0.72\pm 0.10^{***}$ $0.62\pm 0.06^{**}$
+FedBN+ $0.75\pm 0.10^{**}$ $0.63\pm 0.08^{*}$
+FL + PL
+FedAvg $0.74\pm 0.09^{**}$ $0.66\pm 0.09^{*}$
+FedBN+ $0.74\pm 0.10^{**}$ $0.64\pm 0.10^{*}$
+FL + SA
+FedAvg $0.75\pm 0.09^{\text{ns}}$ $0.68\pm 0.10^{\text{ns}}$
+FedBN+ $\mathbf{0.76}\pm 0.10$ $\mathbf{0.70}\pm 0.12$
+PL: Partial Loss, SA: Surgical Aggregation
+(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
+On the NIH test set, surgical aggregation performs comparably to the baseline models with mean AUROC $0.81\pm 0.07$ (all $p>0.05$), apart from the FedAvg PFL baseline ($p\leq 0.001$).
+Furthermore, our method outperforms the existing methods (all $p\leq 0.002$).
+On the CheXpert test set, surgical aggregation outperforms all baselines (all $p\leq 0.03$) and existing methods (all $p\leq 0.004$) with mean AUROC $0.76\pm 0.08$.
+Finally, on the external MIMIC dataset, surgical aggregation outperforms all existing methods with mean AUROC $0.71\pm 0.10$ (all $p\leq 0.001$).
+[…]
+When the mean AUROC scores of the NIH and CheXpert classes is measured separately for the MIMIC test set, we observe that surgical aggregation generalizes better than the baselines and existing methods on NIH classes with mean AUROC $0.76\pm 0.10$ (all $p\leq 0.006$) (4).
+Similarly, on CheXpert classes, our method yields mean AUROC $0.70\pm 0.12$ and outperforms the existing methods (all $p\leq 0.02$), while performing comparably to the baselines (all $p>0.05$), except the CheXpert baseline ($p=0.04$).
+4 Discussion
+Class-heterogeneity is a significant challenge in FL, with most FL methods relying on assumptions that all clients either share the same classes, know the classes of other clients in the network, or have access to a fully annotated public dataset.
+However, there is a significant gap in literature for methods that address class-heterogeneity.
+Our results show that surgical aggregation can train a global model using distributed, class-heterogeneous datasets.
+In the absence of class-heterogeneity, our method performs on par with well-established approaches in literature.
+However, in the presence of class-heterogeneity, our method continues to maintain high-performance and outperforms the baselines.
+It also scales across both IID and non-IID settings.
+On the other hand, existing methods fall short and scale poorly in non-IID settings.
+For example, while FL with partial loss is a close second in performance with IID data, its performance significantly degrades in non-IID settings.
+This is due to feature shift as local updates from a client for non-local classes leads to client drift and performance degradation on non-shared classes (Li et al., 2021; Cohen et al., 2020).
+Unlike our method, FL with partial loss requires clients to account for non-local classes and does not support heterogeneous local classification heads.
+Therefore, selective aggregation is critical to solve class-heterogeneity in non-IID settings.
+Another advantage of our method is its ability to learn a global model that can predict all classes observed in the network.
+While it is popular approach in non-IID settings, PFL entails $K$ clients learning $K$ personalized models.
+This has interesting implications in medical imaging: 1) Each personalized model will yield different predictions for class shared between two clients, thus, raising the question of which model’s prediction should be trusted.
+2) PFL emphasizes local optimization over generalization, thus making personalized models susceptible to poor generalizability (Zhang et al., 2023).
+These advantages enable surgical aggregation to not just facilitate greater collaboration between institutes focusing on different tasks in the same domain, but also enable the countless ”narrowly” focused datasets (e.g., those curated for competitions) to be leveraged to collectively train models with a complete set of abnormalities.
+Despite our results demonstrating the advantages of surgical aggregation, our work has certain limitations: 1) There are no established benchmarks for class-heterogeneity beyond external validation.
+2) Our method requires clients to share the feature extractor, an important consideration for device heterogeneity.
+3) Our experiments were conducted on a simulated FL with a fixed $E=1$ epochs before communication.
+4) Our experiments lacked cross-validation and the initial split of data could influence our method’s performance.
+5) Due to the primary focus on class-heterogeneity, our ablation study did not evaluate the effect of class-imbalance or non-IID data.
+For future work, we intend to explore class-heterogeneity using real-world FL setups (e.g., Flower (Beutel et al., 2020)) and further validate our method across different domains and tasks (such as 3D segmentation).
+Appendix A Convergence Analysis
+[…]
+Definition 1.
+We define class-heterogeneity as the case that covers: 1) Classes shared by all clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \forall i\neq j$, 2) Classes partially shared by some clients in the network, i.e., $\exists c\in C_{i}$ such that $c\in C_{j}\ \exists i\neq j$, and 3) Classes unique to a client in the network, i.e., $\exists c\in C_{i}$ such that $c\not\in C_{j}\ \forall i\neq j$.
+Let us consider that each client $k$ contains private data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$) and is initiated with parameters $\mathbf{w}_{0,k}$ updated using gradient descent.
+We primarily focus this analysis on the convergence of selective aggregation to dynamically build the global classification head. In this case, our distributed optimization problem is defined as:
+$\min_{\mathbf{w}}\left\{(G\circ F)^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}(G\circ F)_{k}^{*}(\mathbf{w})\right\}$ (A.1)
 where the local objective $(G\circ F)_{k}^{*}$ is defined by:
-$(G\circ F)_{k}^{*}\triangleq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$ (2)
-where $L$ is the loss function (e.g., binary cross-entropy) [18].
+$(G\circ F)_{k}^{*}\triangleq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$ (A.2)
+where $L$ is the loss function (e.g., binary cross-entropy) (Li et al., 2019).
 Lemma 1. The optimization of the local objective $(G\circ F)^{*}_{k}$ for client $k$ is controlled by the optimization of $(G\circ F)^{*}_{k}$ on each individual class $c\in C_{k}$.
 Proof: For multi-label classifiers, the loss at client $k$ is computed as the mean loss for each class $c\in C_{k}$ using loss function $L$, such that:
-$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (3)
+$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (A.3)
 where $(y_{k,i})_{c}$ is the ground truth for class $c$ and sample $i\in[n_{k}]$ at client $k$.
@@
   1. 1.
-    Optimizing the global representation block $F$ by locally optimizing client representation blocks $F_{k}$, such that:
-    $\min_{\mathbf{w}}\left\{F^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}^{*}(\mathbf{w})\right\}$ (4)
+    Optimizing the global feature extractor $F$ by locally optimizing client feature extractors $F_{k}$, such that:
+    $\min_{\mathbf{w}}\left\{F^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}^{*}(\mathbf{w})\right\}$ (A.4)
   2. 2.
-    Optimizing the global task block $G$ by locally optimizing client task blocks $G_{k}$, such that:
-    $\min_{\mathbf{w}}\left\{G^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}G_{k}^{*}(\mathbf{w})\right\}$ (5)
+    Optimizing the global classifier $G$ by locally optimizing client classifiers $G_{k}$, such that:
+    $\min_{\mathbf{w}}\left\{G^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}G_{k}^{*}(\mathbf{w})\right\}$ (A.5)
 where the local sub-objectives $F_{k}^{*}=G_{k}^{*}\triangleq(G\circ F)_{k}^{*}$.
 Lemma 2. The optimization of the global sub-objective $F^{*}$ is independent of class-heterogeneity present in the network.
-Proof: Like conventional FL approaches, in our approach the representation block weights $\mathbf{w}_{k,t}^{(F)}$ are always shared with all clients at the end of each communication round, such that:
-$\mathbf{w}_{t,k}^{(F)}=f(\mathbf{w}_{t}^{(F)})$ (6)
-In this case, the convergence of $\mathbf{w}_{k,t}^{(F)}$ is well-established in prior literature [18, 17].
-However, it is not clear whether class-heterogeneity affects the convergence of $\mathbf{w}_{k,t}^{(F)}$.
+Proof: Like conventional FL approaches, in our approach the feature extractor weights $\mathbf{w}_{t,k}^{(F)}$ are always shared with all clients at the end of each communication round, such that:
+$\mathbf{w}_{t}^{(F)}$ $=f(\{\mathbf{w}_{t,k}^{(F)}\ \text{: }\forall k\})$ (A.6)
+$\mathbf{w}_{t,k}^{(F)}$ $=\mathbf{w}_{t}^{(F)}\ \forall k$ (A.7)
+In this case, the convergence of $\mathbf{w}_{t,k}^{(F)}$ is well-established in prior literature (Li et al., 2019, 2021).
+However, it is not clear whether class-heterogeneity affects the convergence of $\mathbf{w}_{t,k}^{(F)}$.
 The optimization of the global sub-objective $F^{*}$ is controlled by the optimization of the local objective $(G\circ F)^{*}_{k}$.
@@
 In other words, the optimization of $F^{*}$ is driven by the optimization of $F_{k}^{*}$, regardless of any class-heterogeneity.
-Since the task block weights are selectively aggregated, we analyze the convergence of the global task block (i.e., $G^{*}$) with respect to each individual class $c\in C$.
+Since the classifier weights are selectively aggregated, we analyze the convergence of the global classifier (i.e., $G^{*}$) with respect to each individual class $c\in C$.
 Lemma 3.
@@
 Proof: Let $K^{\prime}$ be the subset of clients in the network sharing class $c\in C$.
-In our approach, the task-specific weights $(\mathbf{w}_{k,t}^{(G)})_{c}$ for clients $k\in K^{\prime}$ are shared with the clients at the end of each communication round, such that:
-$(\mathbf{w}_{k,t}^{(G)})_{c}=g((\mathbf{w}_{t}^{(G)})_{c})$ (7)
+In our approach, the class-specific weights $(\mathbf{w}_{t,k}^{(G)})_{c}$ for clients $k\in K^{\prime}$ are shared with the clients at the end of each communication round, such that:
+$(\mathbf{w}_{t}^{(G)})_{c}$ $=mean(\{(\mathbf{w}_{t,k}^{(G)})_{c}\ \text{if }c\in C_{k}\ \text{: }\forall k\})$ (A.8)
+$(\mathbf{w}_{t,k}^{(G)})_{c}$ $=(\mathbf{w}_{t}^{(G)})_{c}\ \forall k$ (A.9)
 The global sub-objective $G^{*}$ with respect to the shared class $c$ is controlled by the local optimization of $G_{k}^{*}$ on class $c$ for each clients $k\in K^{\prime}$.
@@
 Using Lemma 2, we conclude that with respect to the class $c$, the optimization of the global objective $(G\circ F)^{*}$ is ensured.
-Putting together Lemmas 3 and 4, we prove that the optimization of global objective $(G\circ F)^{*}$ with respect to class $c\in C$ is always ensured and is derived from previously established convergence analyses, regardless of class-heterogeneity in the network [18, 17].
+Putting together Lemmas 3 and 4, we prove that the optimization of global objective $(G\circ F)^{*}$ with respect to class $c\in C$ is always ensured and is derived from previously established convergence analyses, regardless of class-heterogeneity in the network (Li et al., 2019, 2021).
 Therefore, using Lemma 1, we can confer:
 Theorem 1.
-The optimization of global objective $(G\circ F)^{*}$ is ensured if local objective $(G\circ F)_{k}^{*}$ is optimized for all classes $c\in C_{k}$ using representation block strategy $f$ and task block strategy $g$ and optimization is independent of any class-heterogeneity present in the network.
+The optimization of global objective $(G\circ F)^{*}$ is ensured if local objective $(G\circ F)_{k}^{*}$ is optimized for all classes $c\in C_{k}$ using model aggregation strategy $f$ and the optimization is independent of any class-heterogeneity present in the network.
 To conclude, Theorem 1 says that regardless of the class-heterogeneity present in the network, the convergence of surgical aggregation is guaranteed.
-3.2 Comparison with Baselines
-We compare our method to two baselines, i.e., models trained with well-established approaches in literature that represent an upper bound of the expected model performance, but do not yield a singular global model.
-3.2.1 Individually Trained Baseline
-The individually trained baseline refers to models trained individually on each dataset in a centralized setup using standard deep learning.
-3.2.2 PFL Baseline
-The PFL baseline encapsulates the growing popularity of personalized FL techniques in literature, where each client learns personalized local models in favor of generalized global models [30, 31, 17].
-We implement the PFL baseline such that clients share a representation block, while learning a personalized task block for local classes.
-3.3 Comparison with Existing Methods
-We compare our method to three existing methods for training models by leveraging multiple datasets in aggregate. Unlike the baselines, these methods yield models with a singular task block.
-3.3.1 Centralized
-The centralized method refers to the well-established approach to achieve higher generalizability, where multiple datasets are concatenated locally to train a model in a centralized setup using standard deep learning [8, 7, 6, 9].
-All missing classes in the aggregated dataset are treated as negatives, i.e., $(y_{i})_{c}=0$ for all $c\in C\backslash C_{k}$.
-3.3.2 Vanilla FL
-The vanilla FL method is the conventional approach for FL where datasets are distributed across clients in the network [14, 15, 16].
-Similar to the centralized method, missing classes in each dataset are treated as negatives.
-3.3.3 FL with Partial Loss
-The FL with partial loss method is a modification of the vanilla FL setup, inspired by methods for training multi-label classifiers with missing labels to tackle class-heterogeneity using a modified binary cross-entropy loss [22, 23, 24].
-Jin et al.
-[22] defines the partial loss function for a client $k$ with model parameters $\mathbf{w}$, local classes $C_{k}\subseteq C$, and local training data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples as follows:
-$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (8)
-3.4 Implementation
-3.4.1 Centralized Setup
-We use DenseNet121, pre-trained with ImageNet, to train multi-label classifiers using binary cross-entropy (BCE) loss with batch size of 64.
-Prior to training, task blocks are pre-trained using transfer learning for 15 epochs with learning rate of 5e-3, while keeping representation block weights frozen.
-Then, the model is fine-tuned for 150 epochs with learning rate of 5e-5.
-3.4.2 Federated Setup
-We simulate a FL setup using the above methods with $T=150$ epochs and $E=1$ epochs before communication, resulting in 150 communication rounds.
-Convergence of the global model is determined by the mean of local validation BCE loss for each client (Eq.
-1).
-For each FL approach, we consider the three FL aggregation strategies:
-  • FedAvg, the de-facto FL strategy [13]
-  • FedBN, a variation of FedAvg with local batch normalization [17]. Since FedBN is designed primarily for PFL tasks, it is only included for the PFL baseline model.
-  • FedBN+, a modification of FedBN with pre-trained batch normalization statistics [32].
-3.4.3 Image Preprocessing
-All CXRs are downsampled to $224\times 224$, normalized between 0 and 1, and scaled to ImageNet statistics.
-Random augmentations (rotation, flip, zoom, and contrast) were applied to input images during training.
-3.4.4 Environment
-All models are trained and evaluated using TensorFlow (version 2.8.1) and CUDA (version 12.0) on four NVIDIA RTX A6000 GPUs.
-4 Experimental Setup
-4.1 Datasets
-We use three large-scale CXR datasets that are class- and data-heterogeneous, as shown in Fig. 4.
-Figure 4: Illustration of class- and data-heterogeneity between the NIH, CheXpert, and MIMIC-LT datasets.
-(a) Barplot showing class-heterogeneity between the three datasets as the frequency of occurrence of disease labels.
-The seven shared classes between the NIH and CheXpert datasets are highlighted in bold.
-The MIMIC-LT dataset contains all 20 classes.
-(b) Latent space representation showing data-heterogeneity between all three datasets using t-SNE [33].
-4.1.1 NIH Chest X-Ray 14
-The NIH Chest X-Ray 14 dataset consists of 14 disease labels with $n=112,120$ frontal CXRs from 30,805 patients [5].
-We randomly divide the dataset into training (70%, $n=78,075$), validation (10%, $n=11,079$), and testing (20%, $n=22,966$) splits with no patient leakage.
-4.1.2 CheXpert
-The CheXpert dataset consists of 13 disease labels (seven shared with the NIH dataset) with $n=224,316$ CXRs from 65,240 patients [2].
-All lateral images are discarded to yield $n=191,027$ frontal CXRs.
-Uncertain labels are treated as negatives.
-We randomly divide the dataset into training (70%, $n=133,638$), validation (10%, $n=18,855$), and testing (20%, $n=38,534$) splits with no patient leakage.
-4.1.3 MIMIC-CXR-JPG
-The MIMIC-CXR-JPG dataset is used as our external test set to evaluate the generalizability to unseen data distributions [34, 3].
-The MIMIC dataset consists of 13 disease labels (same as CheXpert) with $n=377,110$ CXRs from 65,379 patients.
-All lateral images are discarded to yield $n=243,324$ frontal CXRs.
-To thoroughly validate our method, we use the CXR-LT expanded set of 26 classes for the MIMIC dataset that provides ground truth for the 20 classes present across the NIH and CheXpert datasets as our fully annotated test set [35, 34].
-Figure 5: Mean AUROC scores for iid simulation experiments across varying levels of class- and data-heterogeneity on the held-out NIH test set (left) and external MIMIC-LT dataset (right).
-(a) Effect of number of clients in the network as a proxy for data-heterogeneity.
-(b) Effect of number of shared classes across all clients in the network as a proxy for class-heterogeneity.
-4.2 Experiments
-4.2.1 Simulation Experiments
-We evaluate the effect of class- and data-heterogeneity using simulated datasets created by sampling iid subsets of the NIH dataset for each client, with varying levels of class- and data-heterogeneity, such that there is no patient leakage and all 14 classes from the NIH dataset appear at least once in each simulated dataset (Constraint 1).
-Since each simulated dataset is an incomplete snapshot of the NIH dataset, we aim to empirically quantify how well our method approximates the performance of the NIH baseline, trained on fully annotated data (Section 3).
-We evaluate our hypothesis across two sub-experiments:
-First, we study the effect of data-heterogeneity across seven runs, each with an increasing number of clients $K\in\{2,3,4,5,6,8,10\}$ as a proxy for data-heterogeneity.
-As the number of clients increases, the sample size of the simulated dataset at each client decreases by factor of $1/K$, thus increasing statistical heterogeneity between clients.
-Second, we study the effect of class-heterogeneity across of seven runs, each with an increasing number of shared classes $\phi\in\{0,2,6,10,12,13,14\}$ as a proxy for class-heterogeneity.
-We keep the number of clients $K=4$ and underlying imaging data at each client constant.
-All models are evaluated on the held-out NIH test set and the external MIMIC-LT dataset.
-For all FL approaches, FedBN+ is used as the aggregation strategy.
-The PFL baseline is not included due to its inability to train singular multi-task models.
-4.2.2 Real-World Experiments
-In addition to evaluating the stability of our framework across varying levels of class- and data-heterogeneity, we evaluate the efficacy of our framework in the real-world using two non-iid large-scale CXR datasets – NIH and CheXpert – to conduct two sub-experiments.
-In the first sub-experiment, we consider the scenario without class-heterogeneity, an ideal scenario for existing FL setups, where only on the seven shared classes between both datasets are included as ground truth annotations (see Fig.
-4).
-In the second experiment, we consider all classes to encapsulate a real-world class- and data-heterogeneous problem by training a model with a complete representation of all 20 classes across both datasets.
-All models are evaluated on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset.
-Figure 6: Average mean AUROC scores for evaluating the effect of data-heterogeneity on the held-out NIH test set and external MIMIC-LT dataset using simulated iid datasets.
-Surgical aggregation is compared with existing methods for each test set using paired t-tests.
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-4.3 Metrics and Statistical Analyses
-Model performance in a multi-label scenario is quantified by the mean of area under the receiver operating characteristic curve (AUROC) scores for each class in the test set.
-If a class in the test set is absent from the set of classes learned by a model, the mean AUROC score is not defined.
-For the simulation experiments, performance between models is compared using paired t-tests between the mean AUROC scores to test for significance.
-For real-world experiments, performance between models is compared using paired t-tests across the granular class-level AUROC scores to test for significance.
-In all cases, statistical significance was defined as $p<0.05$ and the difference between paired samples is normal, indicated by $p\geq 0.05$ using the Shapiro-Wilk test.
-5 Results
-5.1 Simulation Experiments
-Figure 7: Average mean AUROC scores for evaluating the effect of class-heterogeneity on the held-out NIH test set and external MIMIC-LT dataset using simulated iid datasets.
-Surgical aggregation is compared with existing methods for each test set using paired t-tests.
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-5.1.1 Effect of Data-Heterogeneity
-Across varying levels of data-heterogeneity, we observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p<0.005$).
-While FL with partial loss performs comparably to our method, the difference in average mean AUROCs (margin $0.004$) is statistically significant ($p=0.004$).
-On the external MIMIC-LT dataset, our method generalizes better with average mean AUROC $0.69\pm 0.01$, compared to all existing methods (all $p<0.03$).
-Results are detailed in Figs.
-5(a) and 6.
-5.1.2 Effect of Class-Heterogeneity
-Across varying levels of class-heterogeneity, with constant data-heterogeneity, we observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p<0.03$).
-Similarly, the minor differences in average mean AUROCs between FL with partial loss and surgical aggregation are statistically significant (margin $0.004$, $p=0.03$).
-On the external MIMIC-LT dataset, our method generalizes better than all existing methods with average mean AUROC $0.69\pm 0.01$.
-Results are detailed in Figs.
-5(b) and 7.
-Finally, across both sub-experiments, our results indicate that even with varying levels of class- and data-heterogeneity in the network, surgical aggregation provides the closest approximation to the performance of the NIH baseline model (mean AUROC $0.82$, mean margin $0.017$), followed by FL with partial loss as a close second (mean margin $0.021$).
-Table 1: Mean AUROC scores on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset for real-world experiment without class-heterogeneity.
-All values reported as Mean $\pm$ SD.
-Best performing method is highlighted in bold.
-Surgical aggregation is compared with existing methods for each test set using paired t-tests.
-NIH CheXpert MIMIC-LT
-NIH Baseline $0.84\pm 0.05^{\text{ns}}$ $0.70\pm 0.07^{**}$ $0.73\pm 0.10^{**}$
-CheXpert Baseline $0.81\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
-PFL Baseline
- FedAvg $0.83\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
- FedBN $0.83\pm 0.05^{*}$ $0.77\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
- FedBN+ $0.83\pm 0.06^{**}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
-Central $0.82\pm 0.06^{**}$ $0.77\pm 0.07^{\text{ns}}$ $0.76\pm 0.10^{\text{ns}}$
-Vanilla FL
- FedAvg $0.81\pm 0.05^{***}$ $0.74\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
- FedBN+ $0.84\pm 0.05^{*}$ $0.75\pm 0.08^{*}$ $0.75\pm 0.09^{\text{ns}}$
-FL + PartialLoss
- FedAvg $0.82\pm 0.05^{***}$ $0.75\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
- FedBN+ $0.83\pm 0.06^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
-FL + SurgAgg
- FedAvg $0.84\pm 0.05^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
- FedBN+ $\mathbf{0.84}\pm 0.06$ $\mathbf{0.77}\pm 0.08$ $\mathbf{0.76}\pm 0.10$
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-5.2 Real-World Experiments
-5.2.1 Data-Heterogeneous Scenario
-We observe that the purely data-heterogeneous scenario results in similar mean AUROCs across the held-out NIH and CheXpert test sets and the external MIMIC-LT dataset.
-This implies that surgical aggregation is comparable to current FL approaches in the absence of class-heterogeneity (Table 1).
-Specifically, on the NIH test set, surgical aggregation performs comparably to the NIH baseline with mean AUROC $0.84\pm 0.06$ ($p=0.22$), while outperforming other baselines (all $p<0.03$) and existing methods (all $p<0.02$).
-On the CheXpert test set, surgical aggregation performs comparably to the CheXpert baseline model with mean AUROC $0.77\pm 0.08$ ($p=0.51$).
-Our method performs on par with the PFL baselines (all $p>0.05$) and existing methods, while outperforming the NIH baseline ($p=0.002$)
-Finally, on the external MIMIC-LT dataset, we observe that surgical aggregation’s generalizability to unseen data is on par with baselines with mean AUROC $0.76\pm 0.10$ (all $p>0.05$), with the exception of the NIH baseline ($p=0.009$).
-Similarly, our method performs comparably to the existing methods.
-Table 2: Mean AUROC scores on the held-out NIH and CheXpert test sets and external MIMIC-LT dataset for real-world experiment with class-heterogeneity.
-All values reported as Mean $\pm$ SD.
-Best performing method is highlighted in bold.
-Surgical aggregation is compared with existing methods for each test set using paired t-tests.
-NIH CheXpert MIMIC-LT
-NIH Baseline $0.82\pm 0.06^{\text{ns}}$
-CheXpert Baseline - $0.74\pm 0.07^{*}$
-PFL Baseline
- FedAvg $0.79\pm 0.07^{***}$ $0.73\pm 0.07^{***}$
- FedBN $0.81\pm 0.06^{\text{ns}}$ $0.74\pm 0.06^{*}$
- FedBN+ $0.81\pm 0.06^{\text{ns}}$ $0.73\pm 0.07^{*}$
-Central $0.71\pm 0.13^{**}$ $0.72\pm 0.07^{***}$ $0.65\pm 0.10^{***}$
-Vanilla FL
- FedAvg $0.70\pm 0.10^{***}$ $0.68\pm 0.08^{***}$ $0.64\pm 0.09^{***}$
- FedBN+ $0.71\pm 0.12^{***}$ $0.68\pm 0.09^{***}$ $0.66\pm 0.10^{***}$
-FL + PartialLoss
- FedAvg $0.77\pm 0.09^{***}$ $0.68\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
- FedBN+ $0.78\pm 0.07^{***}$ $0.71\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
-FL + SurgAgg
- FedAvg $0.80\pm 0.06^{*}$ $0.73\pm 0.07^{**}$ $0.70\pm 0.09^{\text{ns}}$
- FedBN+ $\mathbf{0.81}\pm 0.07$ $\mathbf{0.76}\pm 0.08$ $\mathbf{0.71}\pm 0.10$
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-5.2.2 Class- and Data-Heterogeneous Scenario
-Table 3: Sub-analysis of mean AUROC scores for shared and unique classes (Fig.
-4) on the held-out NIH and CheXpert test sets for real-world experiment with class-heterogeneity.
-All values reported as Mean $\pm$ SD.
-Best performing method is highlighted in bold.
-Surgical aggregation is compared with existing methods for each test set and sub-group using paired t-tests.
-NIH CheXpert
-Shared Classes Unique Classes Shared Classes Unique Classes
-NIH Baseline $0.84\pm 0.06^{\text{ns}}$ $0.80\pm 0.07^{\text{ns}}$ $0.70\pm 0.08^{**}$
-CheXpert Baseline $0.81\pm 0.07^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.71\pm 0.05^{**}$
-PFL Baseline $0.83\pm 0.06^{**}$ $0.80\pm 0.07^{\text{ns}}$ $0.76\pm 0.07^{\text{ns}}$ $0.70\pm 0.05^{**}$
-Central $0.82\pm 0.05^{*}$ $0.60\pm 0.07^{***}$ $0.74\pm 0.07^{\text{ns}}$ $0.69\pm 0.06^{**}$
-Vanilla FL $0.82\pm 0.06^{**}$ $0.60\pm 0.06^{***}$ $0.73\pm 0.09^{**}$ $0.62\pm 0.05^{**}$
-FL + PartialLoss $0.82\pm 0.05^{*}$ $0.73\pm 0.06^{***}$ $0.74\pm 0.07^{*}$ $0.67\pm 0.04^{**}$
-FL + SurgAgg $\mathbf{0.84}\pm 0.06$ $\mathbf{0.78}\pm 0.07$ $\mathbf{0.77}\pm 0.09$ $\mathbf{0.75}\pm 0.07$
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-Table 4: Sub-analysis of mean AUROC scores for NIH and CheXpert classes on the external MIMIC-LT dataset for real-world experiment with class-heterogeneity.
-All values reported as Mean $\pm$ SD.
-Best performing method is highlighted in bold.
-Surgical aggregation is compared with existing methods for each test set using paired t-tests.
-MIMIC-LT
-NIH Classes CheXpert Classes
-NIH Baseline $**0.73\pm 0.10^{\text{**}}$
-CheXpert Baseline - $0.67\pm 0.11^{*}$
-PFL Baseline
- FedAvg $0.74\pm 0.11^{***}$ $0.69\pm 0.10^{\text{ns}}$
- FedBN $0.70\pm 0.09^{***}$ $0.67\pm 0.10^{\text{ns}}$
- FedBN+ $0.74\pm 0.10^{**}$ $0.68\pm 0.11^{\text{ns}}$
-Central $0.74\pm 0.09^{**}$ $0.66\pm 0.11^{**}$
-Vanilla FL
- FedAvg $0.72\pm 0.10^{***}$ $0.62\pm 0.06^{**}$
- FedBN+ $0.75\pm 0.10^{**}$ $0.63\pm 0.08^{*}$
-FL + PartialLoss
- FedAvg $0.74\pm 0.09^{**}$ $0.66\pm 0.09^{*}$
- FedBN+ $0.74\pm 0.10^{**}$ $0.64\pm 0.10^{*}$
-FL + SurgAgg
- FedAvg $0.75\pm 0.09^{\text{ns}}$ $0.68\pm 0.10^{\text{ns}}$
- FedBN+ $\mathbf{0.76}\pm 0.10$ $\mathbf{0.70}\pm 0.12$
-ns: $p\geq 0.05$, *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$
-We observe that the introduction of class-heterogeneity results in a severe degradation of model performance for the existing methods, while the baseline methods demonstrate high-performance.
-Our results show that surgical aggregation continues to yield high-performance (Table 2).
-Specifically, on the NIH test set, surgical aggregation performs comparably to the baseline models with mean AUROC $0.81\pm 0.07$ (all $p>0.05$), apart from the FedAvg PFL baseline ($p<0.001$).
-Furthermore, our method outperforms the existing methods (all $p<0.002$).
-On the CheXpert test set, surgical aggregation outperforms all baselines (all $p<0.03$) and existing methods (all $p<0.004$) with mean AUROC $0.76\pm 0.08$.
-Finally, on the external MIMIC-LT dataset, surgical aggregation outperforms all existing methods with mean AUROC $0.71\pm 0.10$ (all $p<0.001$).
-We further explore the performance on a granular level by measuring the mean AUROC scores of the shared and the unique classes for the held-out NIH and CheXpert test sets (Table 3).
-We observe that the existing methods perform poorly on the unique classes when compared to the baselines.
-In stark contrast, surgical aggregation demonstrates high-performance on the unique classes.
-On the NIH test set, our method performs on par with the baselines ($p>0.05$), but on the CheXpert test set, surgical aggregation performs comparably to the baselines on the shared classes (both $p>0.05$), while outperforming them on the unique classes (both $p<0.008$).
-Since the mean AUROC scores for baselines are not defined on the expanded MIMIC-LT classes, we explore generalizability by measuring mean AUROC scores of the NIH and CheXpert classes separately on the external MIMIC-LT dataset (Table 4).
-We observe that surgical aggregation generalizes better than the baselines and existing methods on NIH classes with mean AUROC $0.76\pm 0.10$ (all $p<0.006$).
-Similarly, on CheXpert classes, our method yields mean AUROC $0.70\pm 0.12$ and outperforms the existing methods (all $p<0.02$), while performing comparably to the baselines (all $p>0.05$), except the CheXpert baseline ($p=0.04$).
-6 Discussion
-Although the release of countless CXR datasets has spearheaded the utility of deep learning for expert-level detection of diseases in the thorax, the lack of interoperability due to class-heterogeneity has limited their utility beyond training ”narrowly” focused models.
-Therefore, there is a significant gap in literature for methods that address class-heterogeneity and enable training of large-scale models by leveraging non-interoperable datasets in aggregate.
-Our results demonstrate that surgical aggregation enables interoperability between distributed class-heterogeneous datasets to collaboratively develop a global multi-task model.
-In the absence of class-heterogeneity, surgical aggregation performs on par with well-established approaches in literature.
-However, in the presence of class-heterogeneity, our method continues to maintain high-performance and performs on par with, or in some cases outperforms, the baselines.
-More importantly, in the presence of class-heterogeneity, surgical aggregation scales well across both iid and non-iid settings.
-On the other hand, existing methods fall short and scale poorly in non-iid settings.
-For example, while FL with partial loss is a close second in terms of performance with iid data, its performance significantly degrades with non-iid data.
-This is primarily due to feature shift between task-specific weights at each client as local updates from a client for non-local classes leads to client drift and degradation of performance on non-shared classes [17, 6].
-Therefore, the selective aggregation of task-specific weights is critical to tackle class-heterogeneity in non-iid settings and is the crux of our method.
-Another advantage of surgical aggregation is its ability to learn a singular multi-task global model that can predict the complete set of classes observed in the network.
-In stark contrast, the PFL paradigm entails $K$ clients learning $K$ personalized models.
-While it is popular approach to tackle non-iid data, PFL has interesting implications in the clinical environment.
-First, each personalized model will yield different predictions for the same CXR, thus, raising the question of which model’s prediction should be trusted.
-Second, PFL emphasizes local optimization over generalization, thus making personalized models susceptible to poor generalizability on unseen data and shift in labeling schemes [26, 6].
-These advantages enable surgical aggregation to not just facilitate greater collaboration between institutes focusing on different tasks in the same domain, but also enable the countless ”narrowly” focused datasets (e.g., those curated for competitions) to be leveraged to collectively train clinically-useful models with a complete representation of abnormalities.
-Despite our results demonstrating the advantages of surgical aggregation, our work has certain limitations.
-First, surgical aggregation is designed primarily for multi-label classifiers and may not scale for medical image segmentation – a fundamentally different and more difficult problem to solve.
-Second, due to limited literature for class-heterogeneous settings in FL, there are no established benchmarks to compare our method against beyond external validation.
-Third, our method requires all clients to share the same architecture for the representation block, an important consideration for clients with different computational resources.
-Finally, the empirical experiments are conducted on a simulated FL setup – lacking validation in real-world FL networks.
-For future work, we intend to explore solutions for these limitations and further validate our framework for real-world applications.
-7 Conclusion
-Surgical aggregation is a crucial first step towards tackling class-heterogeneity in FL, especially between large-scale CXR datasets.
-Our results show that surgical aggregation can train high-performing models that simultaneously predict the presence of all disease labels present across the datasets and facilitate the development of highly-generalizable models using previously non-interoperable CXR datasets.